### v4

# Coherent manipulation of interacting electron qubits on solid neon

Coherent manipulation of interacting electron qubits on solid neon
Xinhao Li xinhaoli@fas.harvard.edu Department of Physics, Harvard University, Cambridge, Massachusetts 02138, USA Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA    Yizhong Huang Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA Pritzker School of Molecular Engineering, University of Chicago, Chicago, Illinois 60637, USA    Xu Han Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA Pritzker School of Molecular Engineering, University of Chicago, Chicago, Illinois 60637, USA    Xianjing Zhou xianjing.zhou@fsu.edu National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA Department of Mechanical Engineering, FAMU-FSU College of Engineering, Florida State University, Tallahassee, Florida 32310, USA    Amir Yacoby Department of Physics, Harvard University, Cambridge, Massachusetts 02138, USA    Dafei Jin dfjin@nd.edu Department of Physics and Astronomy, University of Notre Dame, Notre Dame, Indiana 46556, USA
(August 24, 2025)
Solid neon has recently emerged as a pristine material host for electron qubits. Single electron-on-solid-neon (eNe) charge qubits have shown extraordinarily long coherence times and high operation fidelities. Realizing two-qubit gates in this platform is the next major step for practical quantum information processing. In this work, we demonstrate frequency- and time-domain coherent manipulation of multiple eNe charge qubits that are coupled by charge-charge interactions. Cross-resonance and bSWAP two-qubit gates are implemented, laying the foundation for universal quantum computing. Inter-qubit coupling strength exceeding 60 MHz is observed, promising fast and high-fidelity two-qubit gates. These results highlight the potential to develop the eNe qubit platform into a compelling quantum computing architecture.
Electrons on solid neon (eNe) is an emerging solid-state qubit platform [1, 2, 3, 4]. Owing to the ultraclean environment provided by the noble-element neon, qubits based on the motional (charge) states of eNe have shown exceedingly long coherence [2]. The inherent electrical addressability of charge states allows for straightforward control and readout by microwave photons in a circuit quantum electrodynamics architecture (cQED) [5, 6], significantly simplifying the circuit design. The small footprint of eNe charge qubits offers another unique advantage for high-volume integration within quantum circuits. To further advance the eNe platform towards a useful quantum information architecture, it is essential to construct systems with multiple interacting qubits and achieve coherent manipulation of all these qubits [7, 4]. In our previous work, we observed two eNe charge qubits trapped around the same designed trapping area and coupled through microwave photons in a common superconducting resonator [2]. Later theoretical and experimental studies show that an imperfect neon surface may play an important role in co-trapping multiple electrons [8, 9, 10]. However, all our observed through-resonator inter-qubit couplings so far are not strong enough for coherent real-time two-qubit gate operations.
In this work, we achieve the first coherent manipulation of multiple interacting electron qubits on solid neon. The rough neon surface provides a natural mechanism to pack several electrons at close distances with strong charge-charge interactions surpassing the qubits’ decoherence rates [11, 4]. In a two-qubit system, spectroscopic measurements show an inter-qubit coupling strength of $J/2\pi=$ 3.35 MHz. By driving one of the two qubits, which is strongly coupled with the superconducting resonator, we are able to demonstrate the cross-resonance and bSWAP types of all-microwave two-qubit gates [12]. The experimental results match well the theoretical simulations based on measured parameters. Furthermore, we observe even stronger coupling in a three-qubit system, with a maximum inter-qubit coupling strength of $J/2\pi=62.5$ MHz. These results lay the foundation for scalable quantum information processing with this emerging platform. The local charge-charge coupling between qubits on solid neon may also provide a unique testbed for quantum simulation of correlated many-electron systems in an ultrapure near-vacuum environment [13, 14].
Multi-qubit systems
Fig. 1. : Schematic of multi-qubit coupling on neon-covered superconducting resonator. (A) Two closely arranged qubits trapped at the open end of a splitting co-planar waveguide superconducting resonator (pink), surrounded by electrical gates (yellow). The blue and gray structures are the intrinsic silicon substrate and ground, respectively. In the experiment, varying gate voltage is symmetrically applied on the left and right trap guards ($\Delta V_{\text{tg}}$) or on the resonator pins ($\Delta V_{\text{res}}$) to tune the qubits’ transition frequency. (B) Schematic of the interactions between a two-qubit system and the superconducting resonator, where a bright qubit $\mathrm{Q_{b}}$ and a dark qubit $\mathrm{Q_{d}}$ are coupled with strength $J$ and $\mathrm{Q_{b}}$ is coupled to the resonator with strength $g_{\mathrm{b}}$. Due to the orthogonal dipole alignment to the field, $\mathrm{Q_{d}}$ is invisible to the resonator. (C) Energy level schematic of the two-qubit system showing possible transitions.
Multiple electron qubits can be trapped in the same trapping area on our device, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonA. They can be driven and read out through microwave photons in the resonator. The electron-photon coupling strength is determined by the magnitude and alignment between the electric dipole moments of the qubits and the local electric field of the microwave photons. For two closely arranged qubits, $\mathrm{Q}_{i}$ and $\mathrm{Q}_{j}$, they can couple through not only microwave photons but also direct charge-charge interactions. The system Hamiltonian can be written as
$\hat{H}_{\mathrm{sys}}=\hat{H}_{\mathrm{r}}+\hat{H}_{\mathrm{q}}+\hat{H}_{\mathrm{r\text{-}q}}+\hat{H}_{\mathrm{q\text{-}q}},$ (1)
where the resonator Hamiltonian $\hat{H}_{\mathrm{r}}$, the qubit Hamiltonian $\hat{H}_{\mathrm{q}}$, the resonator-qubit interaction Hamiltonian $\hat{H}_{\mathrm{r\text{-}q}}$, and the qubit-qubit direct interaction Hamiltonian $\hat{H}_{\mathrm{q\text{-}q}}$ are
$\hat{H}_{\mathrm{r}}=\hbar\omega_{\mathrm{r}}\hat{a}^{\dagger}\hat{a}$, (2)
$\hat{H}_{\mathrm{q}}=\sum_{{i}}\frac{1}{2}\hbar\omega_{i}\hat{\sigma}_{{i}}^{\mathrm{z}},$ (3)
$\hat{H}_{\mathrm{r\text{-}q}}=\sum_{{i}}\hbar g_{i}(\hat{a}^{\dagger}\hat{\sigma}_{i}^{-}+\hat{a}\hat{\sigma}_{{i}}^{+}),$ (4)
and
$\hat{H}_{\mathrm{q\text{-}q}}=\sum_{ij}\hbar J_{ij}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}+\hat{\sigma}_{i}^{-}\hat{\sigma}_{j}^{+})$ (5)
Here, $\omega_{\mathrm{r}}$ is the resonator frequency, $\omega_{i}$ is the transition frequency of $\mathrm{Q}_{i}$, $g_{i}$ is the coupling strength between the resonator and $\mathrm{Q}_{i}$, $\hat{a}$ and $\hat{a}^{\dagger}$ are the annihilation and creation operators of resonator photons, $\hat{\sigma}_{i}^{z}$, $\hat{\sigma}_{i}^{-}$, and $\hat{\sigma}_{i}^{+}$ are the Pauli-$z$, lowering, and raising operators acting on $\mathrm{Q}_{i}$. $J_{ij}$ represents the transverse inter-qubit coupling strength between $\mathrm{Q}_{i}$ and $\mathrm{Q}_{j}$ (ref. [12]). In the following section, we show that the experimentally observed longitudinal (ZZ) coupling strength is small in the system, so we exclude it from the system Hamiltonian model [15, 11].
The imperfect neon surface can pack two electrons at a close distance but with distinct dipole orientations that lead to different electron-photon (qubit-resonator) coupling strengths [16, 5]. As shown in Fig. Coherent manipulation of interacting electron qubits on solid neonB, inside a two-qubit coupled system, one of the qubits with a dipole moment orthogonal to the microwave (MW) electric field of a superconducting resonator can result in a near-zero qubit-resonator coupling strength and negligible dispersive shift of the resonator spectrum. We may call it the “dark” qubit $\mathrm{Q_{d}}$. The other qubit with a dipole moment significantly aligned to the MW electric field has a coupling strength $g_{\mathrm{b}}$. Its quantum states can then be dispersively read out. We call it the “bright” qubit $\mathrm{Q_{b}}$. In such a case, we have a system Hamiltonian with $g_{\mathrm{d}}\sim 0$ and a finite $J$ between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ (ref. [17]).
Fig. 2. : Spectroscopic characterization of coherent two-qubit coupling on solid neon. (A) Two-tone qubit spectroscopy with avoided splittings. The arrows mark the bright qubit $\mathrm{Q_{b}}$ and the dark qubit $\mathrm{Q_{d}}$. The dashed line marks the avoided crossing as in (B). (B) Line cut of the two dressed states at the avoided crossing with splitting magnitude of $J/2\pi=3.35$ MHz. (C) Qubits spectroscopy measured with high-power pulsed readout. Driving pulses with varied frequency ($f_{\text{dr}}$) reach the resonator’s input coupler with power ($P_{\text{dr}}$) increased from -73 dBm to -63 dBm, activating transitions corresponding to the CR and bSWAP two-qubit operations. The measurements were taken when the system was biased at $\Delta V_{\rm tg}=\,$0 V. The phase curves are off-set for visualization.
This two-qubit system can support all-microwave-driven two-qubit operations, including the cross-resonance (CR) and bSWAP types of two-qubit gates, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonC. All-microwave two-qubit gates are widely used for fixed-frequency qubits thanks to their easy implementation, which also preserves qubit coherence by avoiding electrostatic and flux gating [12]. For the CR gate, $\mathrm{Q_{d}}$ is excited by driving $\mathrm{Q_{b}}$ at the frequency of $\mathrm{Q_{d}}$, generating the transition $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$ (refs. [18, 19, 20]). For the bSWAP gate, $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ are driven monochromatically through a two-photon process near the frequency middle point between the two, generating the transition $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$ (refs. [17, 21]).
Two-qubit spectroscopy
We have achieved the strong coupling between the two eNe qubits. Figure Coherent manipulation of interacting electron qubits on solid neonA shows the two-tone qubit spectroscopy measurement. We drive the system with a continuous-wave (CW) tone at a variable frequency, and probe the system with another CW tone at the fixed resonator frequency, with a low power, approximately 130-130 dBm reaching the resonator’s input coupler. Two clear avoided crossings are observed while tuning the bias voltage on the trap guard electrode ($\Delta V_{\rm tg}$). The splitting magnitude of $J/2\pi=$ 3.35 MHz is larger than the linewidth of the hybrid states above and below the splitting, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonB. This indicates the strong coupling between the “bright” qubit ($\mathrm{Q_{b}}$) and “dark” qubit ($\mathrm{Q_{d}}$).
Individually, the bright $\mathrm{Q_{b}}$ has a hyperbolic frequency dependence on $\Delta V_{\rm tg}$, with a charge sweet spot (SS) located near 5.71 GHz, which is about 40 MHz above the resonator frequency at 5.668 GHz. The coupling between $\mathrm{Q_{b}}$ and the resonator dispersively pushes the resonator spectrum, especially when $\mathrm{Q_{b}}$ is tuned to its SS frequency (fig. S0.7). Based on this, we estimate the coupling strength between $\mathrm{Q_{b}}$ and the resonator to be $g_{\mathrm{b}}/2\pi=3.76$ MHz.
In contrast, the interaction between the dark $\mathrm{Q_{d}}$ and the superconducting resonator is much weaker than that of $\mathrm{Q_{b}}$. Meanwhile, the secondary shift of the resonator caused by the dispersive push of $\mathrm{Q_{d}}$ on $\mathrm{Q_{b}}$ is also small. Together, these lead to a negligible $\mathrm{Q_{d}}$-state-dependent shift of the resonator frequency and the absence of visible $\mathrm{Q_{d}}$ spectroscopy in Fig. Coherent manipulation of interacting electron qubits on solid neonA. In addition, the spectroscopic symmetry of the two avoided crossing in against $\Delta V_{\text{tg}}$ indicates that the frequency of the dark $\mathrm{Q_{d}}$ is much less sensitive to $\Delta V_{\rm tg}$ compared to that for the bright $\mathrm{Q_{b}}$, with a nearly constant frequency of 5.726 GHz in the voltage scanning range of Fig. Coherent manipulation of interacting electron qubits on solid neonA. These observations all suggest the misalignment between $\mathrm{Q_{d}}$’s dipole and the microwave and direct current (DC) gating fields.
We attribute the two-qubit coupling to charge-charge interactions at short distances. The spectroscopic observations reveal a coupled two-qubit system represented in Fig. Coherent manipulation of interacting electron qubits on solid neonB, with a finite $g_{\mathrm{b}}$ but negligible $g_{\mathrm{d}}$ in Eq. 4. Due to the distinct coupling strengths of the two qubits individually with the resonator, we can rule out the case of virtual microwave photon exchange between the two qubits through the resonator, in which the inter-qubit exchange strength $J\,=\,g_{\mathrm{b}}g_{\mathrm{d}}(1/\Delta_{\mathrm{b}}+1/\Delta_{\mathrm{d}})/2$, where $\Delta_{\mathrm{b,d}}$ are the detune between the qubits and the resonator [22].
The dark $\mathrm{Q_{d}}$ can be probed by swapping its population to the bright $\mathrm{Q_{b}}$. The swap operation can be realized using the ac-Stark shift of $\mathrm{Q_{b}}$ when the resonator is populated while probing the system with a readout pulse at the resonator frequency. Given the dispersive coupling strength of $\chi/2\pi\approx 0.33$ MHz, the ac-Stark effect will blue shift $\mathrm{Q_{b}}$’s frequency by $2\bar{n}\chi$ (refs. [6, 2]), crossing $\mathrm{Q_{d}}$’s frequency when the averaged intra-cavity photon number $\bar{n}$ is approximately 20. Due to the undetectable coupling between $\mathrm{Q_{d}}$ and the resonator, the probe signal will not affect $\mathrm{Q_{d}}$ directly. With the increase of photon number in the cavity during the readout with sufficient probe power, $\mathrm{Q_{b}}$’s frequency gradually crosses that of $\mathrm{Q_{d}}$, resulting in the states swap and eventually the $\mathrm{Q_{d}}$-state-dependent readout. This effect is turned off when the probe tone is weak, and the ac-Stark shift of $\mathrm{Q_{b}}$ is insufficient to cross the frequency of $\mathrm{Q_{d}}$, as in the case of Fig. Coherent manipulation of interacting electron qubits on solid neonA. We simulate the pulsed readout process in ref. [23], with varied probe power to illustrate the detailed swap process during a 700 ns readout pulse (fig. S0.7), as in the following experiments.
Fig. 3. : Time-domain all-microwave two-qubit operations on solid neon. (A) and (C) Driving pulse length-frequency (A) and driving pulse amplitude-frequency (C) Rabi measurements, following the pulse sequences in the inset, showing the two operations corresponding to cross-resonance (CR, green arrows) and bSWAP (magenta arrows) two-qubit gates. (B) and (D) Simulated $\mathrm{Q_{b}}$ excited state population after the readout pulse in driving pulse length-frequency (B) and driving pulse amplitude-frequency (D) Rabi measurements.
With this method, we perform pulsed qubit spectroscopy when biased at $\mathrm{Q_{b}}$’s SS. A long square pulse (10 µs) with a varying driving frequency $f_{\mathrm{dr}}$ and a driving power ($P_{\mathrm{dr}}$) is first applied to drive the system to a specific state, followed by a short square probe pulse (700 ns) at the resonator frequency $f_{\mathrm{r}}$ with approximately 120-120 dBm power reaching the resonator. The pulse sequence is shown in the inset of Fig. Coherent manipulation of interacting electron qubits on solid neonC. At low driving power, we observe two transitions at 5.711 GHz and 5.726 GHz, corresponding to the frequencies of $\mathrm{Q_{b}}$ at its SS and $\mathrm{Q_{d}}$. The appearance of $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$ transition indicates that the long driving pulse pumps $\mathrm{Q_{d}}$ to a certain population, corresponding to the CR transition of a two-qubit coupled system [20, 18], as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonC. With an increasing driving power, we observe another transition peak ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$) near the middle of the $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ frequencies of 5.7185 GHz. This corresponds to the bSWAP type of two-qubit operation in Fig. Coherent manipulation of interacting electron qubits on solid neonC, involving two-photon excitation [17, 21]. The fact that $\omega_{|1_{\mathrm{b}}1_{\mathrm{d}}\rangle}-\omega_{|0_{\mathrm{b}}1_{\mathrm{d}}\rangle}-\omega_{|1_{\mathrm{b}}0_{\mathrm{d}}\rangle}\simeq 0$ suggests the effective longitudinal (ZZ) interaction in the system is negligible [11, 15]. In summary, through the frequency-domain characterization, we have observed the coherent coupling between two qubits on solid neon, which can support both the CR and the bSWAP types of two-qubit operation.
Two-qubit gate operations
Following the frequency-domain characterization of the coupled system, we perform time-domain two-qubit operations corresponding to the CR and bSWAP gates. Figure Coherent manipulation of interacting electron qubits on solid neonA plots the readout signal following Gaussian-shaped driving pulses with varying lengths and approximately 70-70 dBm reaching the resonator input coupler. The driving pulses are truncated at $\pm 2.5\sigma$ on each side, where $\sigma$ is the standard deviation of the Gaussian envelope. The parameters of the readout pulses are the same as described in the last section. Two oscillation features corresponding to the CR operation near 5.726 GHz and the bSWAP operation near 5.718 GHz can be observed. Based on the system Hamiltonian, we numerically simulate the system evolution through the drive and readout process in the Rabi measurements [23, 24]. Figure Coherent manipulation of interacting electron qubits on solid neonB plots the simulated population of $\mathrm{Q}_{\mathrm{b}}$ after readout. And the evolution of the two-qubit system through the driving process is shown in fig. S0.7. Thanks to the readout process, the excitation of $\mathrm{Q_{d}}$ via the two-qubit operation can be reflected in the $\mathrm{Q_{b}}$ population, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonB, and consequently, the recorded readout signal.
Fig. 4. : Spectroscopic characterization of a three-qubit coupled system on solid neon. (A) Two-tone measured qubits spectroscopy showing the three-qubit coupled system. The dashed line marks the avoided crossing as in (B). (B) The line cut showing large coupling strength between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$ of $J_{\mathrm{12}}/2\pi=$ 62.5 MHz. (C) Calculated eigenstate energy diagram of the three-qubit coupled system, with bare qubit states shown as dashed lines.
The simulated results match well with the experimental observations. For the CR operation (green arrows), $\mathrm{Q_{d}}$ was driven by pumping $\mathrm{Q_{b}}$ at the transition frequency of $\mathrm{Q_{d}}$, whose excited state population reached maximum when the pulse length was approximately 0.8 µs. Meanwhile, the CR operation barely excites $\mathrm{Q_{b}}$, as shown in fig. S0.7. For the bSWAP operation (magenta arrows), the two-photon process drives the two qubits simultaneously, while the oscillation frequency is much lower than the CR operation. Under a square driving pulse with amplitude $A$, and negligible direct driving on the $\mathrm{Q_{d}}$, the oscillation frequency induced by CR operation is $\Omega_{\text{CR}}=AJ/(\Delta_{\mathrm{bd}}+2J^{2}/\Delta_{\mathrm{bd}})$ (ref. [18]). For the bSWAP operation, the oscillation frequency is approximately $\Omega_{\text{bSWAP}}=2A^{2}J/(\Delta_{\mathrm{bd}}+2J^{2}/\Delta_{\mathrm{bd}})^{2}$ (ref. [17]), where $\Delta_{\mathrm{bd}}$ is the frequency separation between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$. When the bSWAP operation performs a $\pi$ rotation, which is about 2 µs in our case, it is locally equivalent to an iSWAP gate with appropriate single-qubit rotations [17].
Furthermore, we performed driving pulse-amplitude-dependent Rabi oscillation measurement with a Gaussian-shaped driving pulse of 800 ns length and output pulse amplitude varying from 0.0 V to 0.25 V, corresponding to zero power to 57-57 dBm reaching the resonator input coupler. Again, we observe oscillating patterns near 5.726 GHz and 5.718 GHz, corresponding to the CR (green arrows) and bSWAP (magenta arrows) operations. We attribute the “waterfall-like” patterns in the frequency range 5.713 through 5.718 GHz in Fig. Coherent manipulation of interacting electron qubits on solid neonC to the effects of the Gaussian pulse shape and the ac-Stark shift of $\mathrm{Q_{b}}$ under higher driving power [23], which are reproduced by simulations in Fig. Coherent manipulation of interacting electron qubits on solid neonD and fig. S0.7.
These results demonstrate, for the first time, time-domain two-qubit operations in a eNe qubit platform, offering the potential for multi-qubit quantum processors based on charge qubits on solid neon. Further optimization of the two-qubit gate and readout is crucial to achieving high-fidelity operations.
Three-qubit strong coupling
We further realized a three-qubit system on solid neon with strong inter-qubit coupling. Figure Coherent manipulation of interacting electron qubits on solid neonA plots the two-tone measured qubit spectrum while varying the gate voltage on the resonator electrode ($\Delta V_{\rm res}$), revealing the energy diagram of this three-qubit system. One of them ($\mathrm{Q}_{1}$) is directly coupled with a superconducting resonator, while the other two ($\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$) are serially coupled to $\mathrm{Q}_{1}$. Compared to $\mathrm{Q}_{1}$, the interactions strengths of $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ with the resonator are weak. We do not observe the vacuum Rabi splitting between them and the resonator. Besides, the dispersive phase shift of the resonator induced by $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ is also small when they are not hybridized with $\mathrm{Q}_{1}$.
In Fig. Coherent manipulation of interacting electron qubits on solid neonB, the line cut around $\Delta V_{\rm res}=0.74$ mV shows a coupling strength of $J_{\textrm{12}}/2\pi=62.5$ MHz between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$. On the other hand, the coupling strength $J_{\textrm{23}}/2\pi$ between $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ is much smaller, at approximately 5 MHz level. The difference in inter-qubit coupling strengths can be caused by the varied distance and dipole orientation between qubits. Besides, since $\mathrm{Q}_{\textrm{2}}$ and $\mathrm{Q}_{\textrm{3}}$ both weakly coupled with the resonator, probably due to unfavorable alignment between their dipole and the microwave field, we attribute the inter-qubit coupling in the three-qubit system to charge-charge interactions instead of virtual photon exchange.
Figure Coherent manipulation of interacting electron qubits on solid neonC shows the calculated eigenstates of the system, which match well with the two-tone spectroscopy. $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$’s SS are both near 6.13 GHz, while $\mathrm{Q}_{3}$’s SS is more detuned to approximately 5.91 GHz. All the three qubits are responsive to the bias voltage on the resonator electrode with various sensitivities. The parameters applied in the calculation is listed in tab. S0.7. Due to the large coupling strength between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$, the state mixing of the three qubits results in the energy diagram of the coupled system, which deviates from the uncoupled qubit spectrum. As a result, the three mixed states are all visible in the two-tone measurement in Fig. Coherent manipulation of interacting electron qubits on solid neonA. These observations indicate the possibilities of simultaneously entangling more than two qubits in the eNe system that may be applied for small- to medium-scale quantum simulations [14, 13].
Discussion and outlook
In this work, we have realized coherent manipulation of multiple electron qubits on solid neon mediated by charge-charge interactions. In a two-qubit system, frequency-domain spectroscopy shows strong inter-qubit coupling. While time-domain measurements demonstrate both cross-resonance and bSWAP two-qubit operations, paving the way for the optimization of high-fidelity two-qubit gates on solid neon. Further, a three-qubit system with over 60 MHz inter-qubit coupling strength shows the potential of building more complex systems with qubits on solid neon.
Charge-charge interactions between qubits at short distances may enable new possibilities for quantum information processing on solid neon. For the two systems studied in this work, qubits can acquire distinct coupling strengths with superconducting resonators. Utilizing such possibilities, qubits with dipole moments oriented normal to the resonator field can serve as quantum memories for nearby coupled qubits [25, 26], or act as a quantum mediator [27]. The variation in qubit-resonator coupling strengths further reflects the complexity of electron trapping on solid neon films and points to the important role of surface profile in determining qubit properties [9, 10, 28, 8].
Establishing coherent charge interactions between adjacent qubits on solid neon is a crucial step in advancing this emerging solid-state qubit platform. Similarly, strong inter-qubit coupling has been demonstrated in semiconductor qubit dots through various interaction mechanisms at short distance [29, 30, 31, 32, 33]. However, a multi-qubit system purely relying on local interactions will pose challenges to integrated control and readout circuits at large qubit volume. Entanglement between distant qubits through superconducting cavity buses could ease these limits posed by local interactions [22, 34, 35, 36]. High-impedance resonators made of high-kinetic inductance thin films [37, 38] and Josephson junction arrays [39, 40] have been applied to achieve coherent interactions between distant semiconductor qubits. Further improvements in the control of individual qubit properties above solid neon with refined film growth, electron trapping and gating mechanisms, as well as the development of a compatible high-impedance resonator [41, 42] would facilitate remote entanglements between distant qubits and hybridize it with other quantum platforms [26, 43]. Looking ahead, the integration of short- and long-distance interactions between qubits on solid neon could pave the way for an on-chip quantum network.
Data availability
The data that support the findings of this study are available from the corresponding authors upon request. Source data are provided with this paper.
Code availability
The codes used to perform the experiments and to analyze the data in this work are available from the corresponding authors upon request.
Acknowledgements.
D.J., A.Y., and X.L. acknowledge support from the Air Force Office of Scientific Research (AFOSR) under Award No. FA9550-23-1-0636 for device fabrication and simulation. D.J. acknowledges support from the Department of Energy (DOE) under Award No. DE-SC0025542 for material growth and characterization. D.J. acknowledges support from the National Science Foundation (NSF) under Award No. OSI-2426768 for theoretical modeling. D.J., X.Z., and Y.H. acknowledge support from the Julian Schwinger Foundation for Physics Research for instrument development. X.H. acknowledges support from the U.S. DOE, Office of Science, Advanced Scientific Computing Research (ASCR) program under Contract No. DE-AC02-06CH11357 as part of the InterQnet quantum networking project. Work performed at the Center for Nanoscale Materials, a U.S. Department of Energy Office of Science User Facility, was supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DEAC02-06CH11357. The authors thank Xuedong Hu and David I. Schuster for helpful discussions.
Supplementary Information for
Coherent manipulation of interacting electron qubits on solid neon
S0.1 Devices and experiments
The data presented in Fig. Coherent manipulation of interacting electron qubits on solid neon and Fig. Coherent manipulation of interacting electron qubits on solid neon were collected on a high-impedance titanium nitride (TiN) splitting superconducting resonator, the same as the one in ref. [9]. The resonator supports a differential mode at 5.668 GHz with a total linewidth of $\kappa/2\pi\,=\,$0.38 MHz. Electrodes with on-chip low-pass filters are connected to the resonator pins and trap guards surrounding the electron-trapping area to apply bias voltages for tuning qubits transition frequency. The device was mounted on a home-made printed circuit board (PCB), sealed in a copper cell with neon filling lines, tungsten filaments, and electrical connectors on top of the cell lid. The cryogenic and room temperature electronics setups are the same as those in ref. [9]. The data presented in Fig. Coherent manipulation of interacting electron qubits on solid neon were collected on a niobium splitting superconducting resonator, the same as the one in refs. [1, 2]. This resonator supports a differential mode at 6.426 GHz, with a total linewidth of $\kappa/2\pi=0.46$ MHz. The sample mounting, cryogenic and room temperature electronic setups were the same as those in refs. [1, 2].
The neon growth procedure is the same as that described in ref. [9]. In summary, the dilution fridge is warmed up to create a temperature gradient from $\sim$27 K to $\sim$25 K between its 4K plate and mixing chamber (MXC) plate. At that moment, neon gas is supplied from a room-temperature gas handling system in the form of puffs and filled onto the device chip in the form of liquid. After a desired amount of neon filling, the fridge heater is turned off to let the fridge cool down to base temperature.
Electrons are deposited on to the chip by sending pulsed current on to the tungsten filaments mounted above the device chip. During the electron firing process, temperature sensed on the MXC plate can increase up to 100 mK. The details of the procedure and discussion of the electron deposition process can be found in ref. [9].
S0.2 Simulation of multi-qubit coupled systems
The populations of qubits in coupled systems are simulated based on their Hamiltonian, using the  QuTiP package [24]. For the numerical data presented in Fig. Coherent manipulation of interacting electron qubits on solid neon, we use Eq. 1 to Eq. 5 as the Hamiltonian, with parameters listed in the following sections and Tab. S0.7. The Lindblad master equation is applied to model the effects of relaxation and dephasing:
$\begin{split}\frac{d\hat{\rho}}{dt}\,=\,&-\frac{i}{\hbar}[\hat{H},\hat{\rho}]\\ &+\frac{\kappa}{2}(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\hat{a}^{\dagger}\hat{a}\hat{\rho}-\hat{\rho}\hat{a}^{\dagger}\hat{a})\\ &+\sum_{i}\frac{\Gamma_{i}}{2}(2\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})\\ &+\sum_{i}\frac{\Gamma_{i}^{\varphi}}{2}(2\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}-(\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})^{2}\hat{\rho}-\hat{\rho}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})^{2}),\end{split}$ (S1)
where $\hat{H}=\hat{H}_{\mathrm{sys}}+\hat{H}_{\mathrm{dr}}$. $\hat{\rho}$ is the system density matrix. $1/\Gamma_{i}$ is the relaxation lifetime $T_{1}$ of $\mathrm{Q}_{i}$, and $2/\Gamma_{i}^{\varphi}$ is the pure dephasing lifetime $T_{\varphi}$ of $\mathrm{Q}_{i}$, which are adopted from relaxation and Ramsey measurements in Fig. S0.7.
To simulate the system under driving pulses starting at time $t=0$ with a time-dependent amplitude $A(t)$ and frequency $\omega_{\mathrm{dr}}$, the driving Hamiltonian can be written as
$\hat{H}_{\mathrm{dr}}\,=\,A(t)\mathrm{cos}(t\omega_{\mathrm{dr}})\sum_{i}\hbar\eta_{i}(\hat{\sigma}_{i}^{+}+\hat{\sigma}_{i}^{-}),$ (S2)
where $\eta_{i}\propto g_{i}$ is the ratio for direct driving on $\mathrm{Q}_{i}$ (ref. [17]). In the two-qubit system discussed in the main text $\eta_{\mathrm{b}}\sim 1$ and $\eta_{\mathrm{d}}\sim 0$. The numerical simulation is performed in the rotating frame at the drive frequency.
To simulate the readout process, we introduce the Hamiltonian of the probing at the resonator ($\hat{H}_{\mathrm{p}}$) to the system Hamiltonian in the rotating frame as:
[…]
where $\epsilon$ is the probing amplitude.
S0.3 Resonator dispersive shift
$\mathrm{Q_{b}}$ of the two-qubit coupled system presented in the main text is coupled to the resonator with its charge sweet-spot (SS) above the resonator frequency. Due to the interaction, the resonator is dispersively shifted when $\mathrm{Q_{b}}$ is biased close to its SS, as shown in Fig. S0.7A. The line-cut of the resonator spectrum when ground-state $\mathrm{Q_{b}}$ is on its SS shows $\chi/2\pi=0.33$ MHz red shift compared to the bare resonator frequency, as shown in Fig. S0.7B. We estimated the coupling strength between $\mathrm{Q_{b}}$ and resonator to be $g_{\mathrm{b}}/2\pi\approx 3.76$ MHz based on this resonator dispersive shift and $\chi\approx g_{\mathrm{b}}^{2}/\Delta_{\mathrm{b}}$ (ref. [6]), where $\Delta_{\mathrm{b}}/2\pi=(\omega_{\mathrm{b}}-\omega_{\mathrm{r}})/2\pi\approx 43$ MHz.
S0.4 Pulsed readout of $\mathrm{Q_{d}}$
We demonstrate the pulsed readout process of $\mathrm{Q_{d}}$ with the following simulation. Initially, the population of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ is set to be 0 and 1, respectively. Then the resonator is populated by sending in a square readout pulse (700 ns) at the resonator frequency with varying amplitude $\epsilon$. The evolution of the coupled system is calculated with the system Hamiltonian Eq. 1 and the probing Hamiltonian Eq. S3, considering relaxation and dephasing with Eq. S1.
Figure S0.7 shows the evolution of intra-cavity photon number $\bar{n}$, $\mathrm{Q_{b}}$, and $\mathrm{Q_{d}}$ excited state population under varying driving amplitude. When the probe power is small (Fig. S0.7A and Fig.S0.7B), the ac-Stark shift of $\mathrm{Q_{b}}$ caused by the probe pulse is also small, $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ weakly exchange population at high frequency approximately $\sqrt{\Delta_{\mathrm{bd}}^{2}+4J^{2}}$, where $J$ is the inter-qubit coupling strength and $\Delta_{\mathrm{bd}}$ are the qubits detune. With increased photon number, the blue ac-Stark shift $2\chi\bar{n}$ causes the $\mathrm{Q_{b}}$ across $\mathrm{Q_{d}}$ frequency, resulting in their population swap when $\bar{n}$ approaches 20, as shown in Fig. S0.7C. Further increasing the probing amplitude could populate the resonator faster and make the population swap happen even sooner. These simulation results reveal the readout process we applied to probe $\mathrm{Q_{d}}$, whose direct interaction strength with the resonator is negligible. We chose the 700 ns readout length to balance signal strength with the fast decaying of $\mathrm{Q_{b}}$, as illustrated in the qubit decoherence section.
S0.5 Qubits decoherence
We evaluated $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$’s coherence via relaxation and Ramsey measurements when $\mathrm{Q_{b}}$ was biased at its SS. $\mathrm{Q_{b}}$ was directly driven by sending pumps through the resonator at its frequency. In contrast, $\mathrm{Q_{d}}$ was driven by pumping $\mathrm{Q_{b}}$ at $\mathrm{Q_{d}}$’s frequency, corresponding to the cross-resonance two-qubit operation. $\mathrm{Q_{d}}$ was probed with the pulsed readout method described in the main text and the previous sections. Figure S0.7A and S0.7B show the relaxation of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ with $T_{1}$ of 1.8 µs and 30.3 µs, respectively, fitted from the measured in-phase signal. Figure S0.7C shows the Ramsey fringes of $\mathrm{Q_{b}}$, with a $T_{2}^{*}$ of 2.56 µs fitted from the in-phase signal. The Ramsey measurement result of $\mathrm{Q_{d}}$ is more complicated, as we observed multiple frequency components of the measured in-phase signal, as shown in Fig. S0.7D and  S0.7E. Numerical simulation of the Ramsey measurement process with a stable single frequency of $\mathrm{Q_{d}}$ reveals that the population swap between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ induced by the inter-qubit coupling will not affect the final probed state and measured signal. Therefore, we attribute the beating features in Fig. S0.7D to fluctuations of the $\mathrm{Q_{d}}$’s frequency during data acquisition, which has also occurred on other qubits measured on the same device, as presented in ref. [9]. We fitted the data in Fig. S0.7D with a three-frequency model, which resulted in a $T_{2}^{*}$ of 32.99 µs. The measured relaxation and decoherence rates are then used in the numerical simulation of the system, as shown in Tab. S0.7.
S0.6 Numerical simulation parameters
Table S0.7 summarizes the simulation parameters to build the Hamiltonian of the two-qubit coupled system and to account for the decoherence. Since the longitudinal interaction is small, as discussed in the main text, ZZ coupling is excluded in the simulation. The resonator has a frequency of 5.668 GHz with linewidth $\kappa/2\pi=0.38$ MHz. The interaction strength between the bright qubit and the resonator is $g_{\mathrm{b}}/2\pi=3.76$ MHz. In addition, the driving amplitude $A$ in MHz used in the simulation is scaled with $A=735\times\mathcal{A}$, where $\mathcal{A}$ is the experimental driving pulse amplitude in $V$. The readout process is simulated with 700 ns square pulse with probing amplitude $\epsilon$ at 18 MHz.
Figure S0.7 shows the simulated qubit excited state population before the readout process in the Rabi measurements, indicated as the arrow time stamp in the inset. The oscillation features at 5.726 GHz and 5.718 GHz, corresponding to the cross-resonance (CR) and bSWAP operation, match well with the experimental result in Fig. Coherent manipulation of interacting electron qubits on solid neon of the main text. It confirms that the observed oscillation features are the result of the two-qubit operations instead of the probing pulse following that. In Fig. S0.7A and S0.7C, when $\mathrm{Q_{d}}$ is excited by driving $\mathrm{Q_{b}}$ at 5.726 GHz, the $\mathrm{Q_{b}}$’s population barely changes. After the readout process, the $\mathrm{Q_{d}}$’s population is swapped onto $\mathrm{Q_{b}}$, as shown in Fig. Coherent manipulation of interacting electron qubits on solid neonB and Coherent manipulation of interacting electron qubits on solid neonD of the main text.
To calculate the eigenstates of the three-qubit coupled system, as in the main text Fig. Coherent manipulation of interacting electron qubits on solid neon, we modeled qubits 1, 2, and 3 with hyperbolic frequency dependency on the gate voltage $\Delta V_{\rm res}$ as: $f_{i}=\sqrt{(\alpha_{i})^{2}+(\beta_{i}\times(\Delta V_{\rm res}-\delta_{i}))^{2}}$. In Tab. S0.7, we list the parameters used to generate Fig. Coherent manipulation of interacting electron qubits on solid neonC in the main text. In addition, we used $J_{\textrm{12}}/2\pi\,=62.5$ MHz and $J_{\textrm{23}}/2\pi\,=5.0$ MHz. We also omit longitudinal interactions between qubits in this calculation.
S0.7 “Waterfall” features in pulse amplitude-dependent Rabi oscillation
The “waterfall-like” pattern measured in pulse amplitude-dependent Rabi oscillation is also replicated in the numerical simulation, by realistically modeling the driving pulse with the same Gaussian envelope used in experiments. We attribute it to the combined effects of the pulse shape and ac-Stark shift of $\mathrm{Q_{b}}$ under driving at the frequency middle point between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ with high power. Figure S0.7 plots the simulation when we replace the Gaussian-shaped driving pulse with square pulses. Under such conditions, the “waterfall-like” pattern vanishes while we could still observe the red shift of the $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$ transition at higher power, caused by the ac-Stark shift of $\mathrm{Q_{b}}$.
Fig. S1. : Resonator dispersive shift caused by $\mathrm{Q_{b}}$. (A) Resonator spectrum probe with varying bias on trap guard, showing the dispersive shift when $\mathrm{Q_{b}}$ is close to resonator frequency. (B) Transmission line cuts at bare resonator (blue dots and curve) and when $\mathrm{Q_{b}}$ is biased at its sweet-spot (orange dots and curve).
Fig. S2. : Simulation of intra-cavity photon number and qubit population evolution with varying probing amplitude. 700 ns evolution simulated with probing amplitude $\epsilon$ of (A) 5.6 MHz (-130 dBm reaching the resonator), (B) 10 MHz (-125 dBm reaching the resonator) and (C) 18 MHz (-120 dBm reaching the resonator), respectively. Black curves show intra-cavity photon number $\bar{n}$. Blue curves show $\mathrm{Q_{b}}$ population and orange curves show $\mathrm{Q_{d}}$ population.
Fig. S3. : Relaxation and decoherence of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$. (A) and (B) Relaxation measurements of $\mathrm{Q_{b}}$ (A) and $\mathrm{Q_{d}}$ (B) with fitted $T_{1}$ of 1.8 µs and 30.3 µs, respectively. (C) and (D) Ramsey measurements of $\mathrm{Q_{b}}$ (C) and $\mathrm{Q_{d}}$ (D) with fitted $T_{2}^{*}$ of 2.56 µs and 32.99 µs, respectively. (E) Frequency components of the Ramsey fringes in (D), revealing frequency fluctuations of $\mathrm{Q_{d}}$ during the measurement. (F) Numerical simulation of Ramsey measurement of $\mathrm{Q_{d}}$, indicating that the beating pattern in (C) is not intrinsically caused by the coupled system.
Fig. S4. : Numerical simulation of qubit excited state population before the readout process in Rabi measurements. (A) and (B) Simulation results of $\mathrm{Q_{b}}$ (A) and $\mathrm{Q_{d}}$ (B) population after driving pulse in the pulse length-frequency Rabi measurements. (C) and (D) Simulation results of $\mathrm{Q_{b}}$ (C) and $\mathrm{Q_{d}}$ (D) population after driving pulse in the pulse amplitude-frequency Rabi measurements.
Fig. S5. : Numerical simulation of pulse amplitude-dependent Rabi oscillation with step driving pulses. Simulation results of $\mathrm{Q_{b}}$ (A) and $\mathrm{Q_{d}}$ (B) population evolution, in which the “waterfall-like” pattern observed in experiments vanishes.
Tab. S1. : Summary of the qubit parameters used for simulating the two-qubit coupled system as in Fig. Coherent manipulation of interacting electron qubits on solid neon and Fig. S0.7.
$\omega_{i}/2\pi$ (GHz) $\Gamma_{i}/2\pi$ (MHz) $\Gamma_{i}^{\varphi}/2\pi$ (MHz) $g_{i}/2\pi$ (MHz) $J/2\pi$ (MHz)
Bright $\mathrm{Q_{b}}$ 5.7112 0.088 0.036 3.76 3.35
Dark $\mathrm{Q_{d}}$ 5.7255 0.0053 0.0044 0.0
Tab. S2. : Summary of the qubit parameters used for modeling the three-qubit coupled system as in Fig. Coherent manipulation of interacting electron qubits on solid neonC of the main text.
$\alpha_{i}$ (GHz) $\beta_{i}$ (GHz/mV) $\delta_{i}$ (mV)
Qubit 1 6.130 1.325 0.0
Qubit 2 6.135 0.218 -4.0
Qubit 3 5.915 1.30 0.08

### v5

@@
 Xinhao Li xinhaoli@fas.harvard.edu Department of Physics, Harvard University, Cambridge, Massachusetts 02138, USA Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA    Yizhong Huang Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA Pritzker School of Molecular Engineering, University of Chicago, Chicago, Illinois 60637, USA    Xu Han Center for Nanoscale Materials, Argonne National Laboratory, Lemont, Illinois 60439, USA Pritzker School of Molecular Engineering, University of Chicago, Chicago, Illinois 60637, USA    Xianjing Zhou xianjing.zhou@fsu.edu National High Magnetic Field Laboratory, Tallahassee, Florida 32310, USA Department of Mechanical Engineering, FAMU-FSU College of Engineering, Florida State University, Tallahassee, Florida 32310, USA    Amir Yacoby Department of Physics, Harvard University, Cambridge, Massachusetts 02138, USA    Dafei Jin dfjin@nd.edu Department of Physics and Astronomy, University of Notre Dame, Notre Dame, Indiana 46556, USA
-(August 24, 2025)
-Solid neon has recently emerged as a pristine material host for electron qubits.
-Single electron-on-solid-neon (eNe) charge qubits have shown extraordinarily long coherence times and high operation fidelities.
-Realizing two-qubit gates in this platform is the next major step for practical quantum information processing.
-In this work, we demonstrate frequency- and time-domain coherent manipulation of multiple eNe charge qubits that are coupled by charge-charge interactions.
-Cross-resonance and bSWAP two-qubit gates are implemented, laying the foundation for universal quantum computing.
-Inter-qubit coupling strength exceeding 60 MHz is observed, promising fast and high-fidelity two-qubit gates.
-These results highlight the potential to develop the eNe qubit platform into a compelling quantum computing architecture.
-Electrons on solid neon (eNe) is an emerging solid-state qubit platform [1, 2, 3, 4].
-Owing to the ultraclean environment provided by the noble-element neon, qubits based on the motional (charge) states of eNe have shown exceedingly long coherence [2].
-The inherent electrical addressability of charge states allows for straightforward control and readout by microwave photons in a circuit quantum electrodynamics architecture (cQED) [5, 6], significantly simplifying the circuit design.
-The small footprint of eNe charge qubits offers another unique advantage for high-volume integration within quantum circuits.
-To further advance the eNe platform towards a useful quantum information architecture, it is essential to construct systems with multiple interacting qubits and achieve coherent manipulation of all these qubits [7, 4].
-In our previous work, we observed two eNe charge qubits trapped around the same designed trapping area and coupled through microwave photons in a common superconducting resonator [2].
-Later theoretical and experimental studies show that an imperfect neon surface may play an important role in co-trapping multiple electrons [8, 9, 10].
-However, all our observed through-resonator inter-qubit couplings so far are not strong enough for coherent real-time two-qubit gate operations.
-In this work, we achieve the first coherent manipulation of multiple interacting electron qubits on solid neon.
-The rough neon surface provides a natural mechanism to pack several electrons at close distances with strong charge-charge interactions surpassing the qubits’ decoherence rates [11, 4].
-In a two-qubit system, spectroscopic measurements show an inter-qubit coupling strength of $J/2\pi=$ 3.35 MHz.
-By driving one of the two qubits, which is strongly coupled with the superconducting resonator, we are able to demonstrate the cross-resonance and bSWAP types of all-microwave two-qubit gates [12].
-The experimental results match well the theoretical simulations based on measured parameters.
-Furthermore, we observe even stronger coupling in a three-qubit system, with a maximum inter-qubit coupling strength of $J/2\pi=62.5$ MHz.
-These results lay the foundation for scalable quantum information processing with this emerging platform.
-The local charge-charge coupling between qubits on solid neon may also provide a unique testbed for quantum simulation of correlated many-electron systems in an ultrapure near-vacuum environment [13, 14].
-Multi-qubit systems
+Electrons trapped on solid neon surfaces serve as low-noise charge qubits with long coherence times and high operational fidelities.
+Such charge qubits offer full electrical control and compact device footprints, convenient for scaling up with quantum circuits.
+Realizing two-qubit gates on this platform is a critical step towards practical quantum information processing.
+In this work, we report the first experimental demonstration of coherent manipulation of multiple interacting electron-on-solid-neon (eNe) charge qubits.
+By exploiting the electrons naturally confined in close proximity by the surface structures of solid neon, we have achieved a direct qubit-qubit coupling strength of up to 62.5 MHz, as well as implemented cross-resonance (CR) and bSWAP two-qubit gates using global microwave drives.
+The natural electron confinement by solid neon mitigates the high-density-wiring challenge, simplifies the multi-qubit control, and establishes a unique path to scale up the eNe qubit platform.
+Scalable quantum processors require the constituent qubits to have long coherence, fast and accurate gates, and high-connectivity integration.
+To that extent, electron-based qubits have unique advantages for large-volume solid-state system scaling, thanks to electrons’ small footprints.
+For example, spin qubits based on electrons in gate-defined quantum dots or implanted dopants exhibit very long coherence, and form multi-qubit arrays [burkard2023semiconductor, Petta2005coherent, borsoi2024shared, vandersypen2017interfacing, zwerver2022qubits, hu2025single].
+However, this scaling requires complex gate stacks for electron confinement and spin detection at the individual qubit scale.
+Meanwhile, the necessity for global or local magnetic field gradients to address electron spin adds another layer of hardware complexity for integration.
+Those overheads impede the volume growth of electron-based qubit arrays [vandersypen2017interfacing, borsoi2024shared].
+Electron-on-solid-neon (eNe) offers an emerging solid-state qubit platform that combines excellent coherence and minimal device overhead [zhou2022single, zhou2024electron, guo2024quantum, jennings2024quantum].
+In this system, single electrons sit in a near-vacuum environment a few nanometers above the neon surface, where the noble-gas substrate provides superior noise isolation [li2025noise].
+Qubits based on the charge (motional) states of eNe have shown exceedingly long coherence [zhou2024electron].
+The inherent electrical addressability of charge states allows for straightforward and high-fidelity single-qubit gates with microwave photons in a circuit quantum electrodynamics (cQED) architecture [schuster2010proposal, blais2021circuit], significantly reducing the hardware complexity compared to electron spin qubits.
+To further advance the eNe platform towards a functional quantum information architecture, it is essential to construct systems with multiple interacting qubits and realize their coherent manipulation [platzman1999quantum, jennings2024quantum].
+Previously, we observed two eNe charge qubits co-trapped on the same device and simultaneously coupled to a common superconducting resonator [zhou2024electron].
+However, the through-resonator inter-qubit coupling was not strong enough for coherent real-time two-qubit gate operations.
+Nonetheless, recent theoretical and experimental studies suggest that solid-neon surface morphology offers a natural mechanism to trap multiple electrons at close distances [kanai2024single, li2025noise, zheng2025surface].
+It shows the promise of creating strong interactions between adjacent qubits maintained solely by the neon surface.
+This can greatly simplify the originally complex electrodes in typical electron-based qubit systems [vandersypen2017interfacing] and reduce gate-introduced electrical noise.
 Fig.
 1.
-: Schematic of multi-qubit coupling on neon-covered superconducting resonator.
-(A) Two closely arranged qubits trapped at the open end of a splitting co-planar waveguide superconducting resonator (pink), surrounded by electrical gates (yellow).
-The blue and gray structures are the intrinsic silicon substrate and ground, respectively.
-In the experiment, varying gate voltage is symmetrically applied on the left and right trap guards ($\Delta V_{\text{tg}}$) or on the resonator pins ($\Delta V_{\text{res}}$) to tune the qubits’ transition frequency.
-(B) Schematic of the interactions between a two-qubit system and the superconducting resonator, where a bright qubit $\mathrm{Q_{b}}$ and a dark qubit $\mathrm{Q_{d}}$ are coupled with strength $J$ and $\mathrm{Q_{b}}$ is coupled to the resonator with strength $g_{\mathrm{b}}$.
-Due to the orthogonal dipole alignment to the field, $\mathrm{Q_{d}}$ is invisible to the resonator.
-(C) Energy level schematic of the two-qubit system showing possible transitions.
-Multiple electron qubits can be trapped in the same trapping area on our device, as shown in Fig.
-Coherent manipulation of interacting electron qubits on solid neonA.
-They can be driven and read out through microwave photons in the resonator.
-The electron-photon coupling strength is determined by the magnitude and alignment between the electric dipole moments of the qubits and the local electric field of the microwave photons.
-For two closely arranged qubits, $\mathrm{Q}_{i}$ and $\mathrm{Q}_{j}$, they can couple through not only microwave photons but also direct charge-charge interactions.
-The system Hamiltonian can be written as
-$\hat{H}_{\mathrm{sys}}=\hat{H}_{\mathrm{r}}+\hat{H}_{\mathrm{q}}+\hat{H}_{\mathrm{r\text{-}q}}+\hat{H}_{\mathrm{q\text{-}q}},$ (1)
-where the resonator Hamiltonian $\hat{H}_{\mathrm{r}}$, the qubit Hamiltonian $\hat{H}_{\mathrm{q}}$, the resonator-qubit interaction Hamiltonian $\hat{H}_{\mathrm{r\text{-}q}}$, and the qubit-qubit direct interaction Hamiltonian $\hat{H}_{\mathrm{q\text{-}q}}$ are
-$\hat{H}_{\mathrm{r}}=\hbar\omega_{\mathrm{r}}\hat{a}^{\dagger}\hat{a}$, (2)
-$\hat{H}_{\mathrm{q}}=\sum_{{i}}\frac{1}{2}\hbar\omega_{i}\hat{\sigma}_{{i}}^{\mathrm{z}},$ (3)
-$\hat{H}_{\mathrm{r\text{-}q}}=\sum_{{i}}\hbar g_{i}(\hat{a}^{\dagger}\hat{\sigma}_{i}^{-}+\hat{a}\hat{\sigma}_{{i}}^{+}),$ (4)
-and
-$\hat{H}_{\mathrm{q\text{-}q}}=\sum_{ij}\hbar J_{ij}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}+\hat{\sigma}_{i}^{-}\hat{\sigma}_{j}^{+})$ (5)
-Here, $\omega_{\mathrm{r}}$ is the resonator frequency, $\omega_{i}$ is the transition frequency of $\mathrm{Q}_{i}$, $g_{i}$ is the coupling strength between the resonator and $\mathrm{Q}_{i}$, $\hat{a}$ and $\hat{a}^{\dagger}$ are the annihilation and creation operators of resonator photons, $\hat{\sigma}_{i}^{z}$, $\hat{\sigma}_{i}^{-}$, and $\hat{\sigma}_{i}^{+}$ are the Pauli-$z$, lowering, and raising operators acting on $\mathrm{Q}_{i}$.
-$J_{ij}$ represents the transverse inter-qubit coupling strength between $\mathrm{Q}_{i}$ and $\mathrm{Q}_{j}$ (ref.
-[12]).
-In the following section, we show that the experimentally observed longitudinal (ZZ) coupling strength is small in the system, so we exclude it from the system Hamiltonian model [15, 11].
-The imperfect neon surface can pack two electrons at a close distance but with distinct dipole orientations that lead to different electron-photon (qubit-resonator) coupling strengths [16, 5].
-As shown in Fig.
-Coherent manipulation of interacting electron qubits on solid neonB, inside a two-qubit coupled system, one of the qubits with a dipole moment orthogonal to the microwave (MW) electric field of a superconducting resonator can result in a near-zero qubit-resonator coupling strength and negligible dispersive shift of the resonator spectrum.
-We may call it the “dark” qubit $\mathrm{Q_{d}}$.
-The other qubit with a dipole moment significantly aligned to the MW electric field has a coupling strength $g_{\mathrm{b}}$.
-Its quantum states can then be dispersively read out.
-We call it the “bright” qubit $\mathrm{Q_{b}}$.
-In such a case, we have a system Hamiltonian with $g_{\mathrm{d}}\sim 0$ and a finite $J$ between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ (ref.
-[17]).
+: Schematic of multi-qubit coupling on solid neon.
+a, Solid neon creates multi-qubit host spots and supports inter-qubit coupling at a short distance.
+b, Schematic of the interactions between a two-qubit system on solid neon and a superconducting resonator, where a bright qubit $\mathrm{Q_{b}}$ and a dark qubit $\mathrm{Q_{d}}$ are coupled with strength $J$, and $\mathrm{Q_{b}}$ is coupled to the resonator with strength $g_{\mathrm{b}}$.
+Due to the orthogonal dipole alignment to the field, as marked by the black arrow, $\mathrm{Q_{d}}$ is invisible to the resonator.
+c, Energy-level schematic of the two-qubit system in b, illustrating the possible transitions.
+Both the cross-resonance (CR) and the bSWAP type two-qubit operations are shown, where the $\mathrm{Q_{d}}$ are driven by microwave gates mediated through the $\mathrm{Q_{b}}$.
+In this work, we explore this novel potential offered by solid neon and achieve the first coherent manipulation of multiple interacting electron qubits on this platform.
+We utilize the neon surface structure as a natural mechanism to pack several electrons at close distances to establish strong charge-charge interactions [beysengulov2024coulomb, jennings2024quantum].
+Further, we prove that the natural confinement of electrons by solid neon offers a unique engineering resource to qubit-resonator and inter-qubit coupling, enabling multi-qubit operation with simplified wiring.
+In a two-qubit system, spectroscopic measurements show an inter-qubit coupling strength of $J/2\pi$ = 3.35 MHz, surpassing the qubits’ decoherence rates.
+By driving one of the two qubits, which is the only one strongly coupled with the superconducting resonator, we demonstrate the cross-resonance and bSWAP types of all-microwave two-qubit operations [krantz2019quantum], with just a single direct-current (DC) qubit bias.
+The experimental results match well with the theoretical simulations based on the measured parameters.
+Furthermore, we observe even stronger coupling in a three-qubit system, with a maximum qubit interaction strength of $J/2\pi$ = 62.5 MHz.
+Together with previously achieved long coherence and noise isolation, this result on coherent multi-qubit manipulation completes the eNe platform with crucial components to overcome key scaling challenges faced by electron-based qubits.
+Multi-qubits on solid neon
+The solid neon film can support the trapping of multiple electrons within a compact area without introducing complex electrical gate stacks at the individual qubit scale.
+Experimental and theoretical investigations show that surface structures at the 10-nm scale facilitate electron trapping on solid neon [zheng2025surface, kanai2024single].
+Those surface structures can originate from the transfer of the underneath substrate roughness or the solidification process during the neon layer growth [leiderer2025surface].
+As illustrated in Fig.
+Coherent manipulation of interacting electron qubits on solid neona, the solid neon film naturally supports closely spaced electron traps and enables local qubit coupling through strong charge-charge interactions.
+The trapping of electrons on solid neon films also creates the possibilities for interaction engineering with simplified circuit wiring.
+In the cQED architecture, electrons can be driven and read out through microwave photons in the resonator.
+The electron-photon coupling strength is determined by the magnitude of the qubit’s dipole moment and its alignment with the local microwave (MW) electric field.
+The neon surface morphology can lead to distinct electron-photon couplings [koolstra2019coupling, schuster2010proposal].
+As depicted in Fig.
+Coherent manipulation of interacting electron qubits on solid neonb, inside a closely spaced two-qubit coupled system, one of the qubits with a dipole orthogonal to the MW field of a superconducting resonator exhibits a near-zero qubit-resonator coupling strength and negligible resonator dispersive shift.
+We refer to this qubit as the “dark” qubit $\mathrm{Q_{d}}$.
+The other qubit, which has a dipole significantly aligned to the MW field, is coupled to the resonator with a strength of $g_{\mathrm{b}}$.
+This qubit is referred to as the “bright” qubit $\mathrm{Q_{b}}$ and its quantum states can be read out dispersively.
+The charge-charge interaction between the two qubits results in the inter-qubit coupling strength $J$.
 Fig.
@@
 : Spectroscopic characterization of coherent two-qubit coupling on solid neon.
-(A) Two-tone qubit spectroscopy with avoided splittings.
-The arrows mark the bright qubit $\mathrm{Q_{b}}$ and the dark qubit $\mathrm{Q_{d}}$.
-The dashed line marks the avoided crossing as in (B).
-(B) Line cut of the two dressed states at the avoided crossing with splitting magnitude of $J/2\pi=3.35$ MHz.
-(C) Qubits spectroscopy measured with high-power pulsed readout.
-Driving pulses with varied frequency ($f_{\text{dr}}$) reach the resonator’s input coupler with power ($P_{\text{dr}}$) increased from -73 dBm to -63 dBm, activating transitions corresponding to the CR and bSWAP two-qubit operations.
+a, Continuous-wave (CW) two-tone qubit spectroscopy with avoided crossings on the resonator’s phase response.
+The dashed lines indicate the bare transition frequencies of the dark qubit $\mathrm{Q_{d}}$ and the bright qubit $\mathrm{Q_{b}}$.
+b, The line cut at the avoided crossing shows a splitting corresponding to a charge-charge interaction strength of $J/2\pi=3.35$ MHz, which exceeds the hybrid linewidth $(\gamma_{\textrm{b}}+\gamma_{\textrm{d}})/4\pi=0.56$ MHz.
+c, Qubits spectroscopy measured with high-power pulsed readout, in which $\mathrm{Q_{b}}$ served as the mediator between $\mathrm{Q_{d}}$ and the resonator.
+The inset shows the pulse sequence: a long square pulse (10 µs) with variable frequency $f_{\mathrm{dr}}$ and power $P_{\mathrm{dr}}$ drive the system to a specific state, followed by a short square probe pulse (0.7 µs) at the resonator frequency $f_{\mathrm{r}}$ with approximately 120-120 dBm power reaching the resonator.
+The readout pulse induces the frequency crossing of the two qubits and the swap of their populations, revealing the state information of the dark qubit $\mathrm{Q_{d}}$.
+Drive pulse power $P_{\text{dr}}$ reaching the resonator’s input coupler increases from 73-73 dBm to 63-63 dBm, activating transitions corresponding to the CR ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$) and bSWAP ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$) two-qubit operations.
 The measurements were taken when the system was biased at $\Delta V_{\rm tg}=\,$0 V.
 The phase curves are off-set for visualization.
-This two-qubit system can support all-microwave-driven two-qubit operations, including the cross-resonance (CR) and bSWAP types of two-qubit gates, as shown in Fig.
+The structurally induced dipole orientations distribution enables coherent multi-qubit operations to be driven with only global control, reducing the complexity of the local gate stacks.
+The two-qubit system described above supports all-microwave-driven two-qubit operations, including cross-resonance (CR) and bSWAP gates, as presented in Fig.
 Coherent manipulation of interacting electron qubits on solid neonc.
-All-microwave two-qubit gates are widely used for fixed-frequency qubits thanks to their easy implementation, which also preserves qubit coherence by avoiding electrostatic and flux gating [12].
+All-microwave gates are widely used for fixed-frequency qubits thanks to their easy implementation, which also preserves qubit coherence by avoiding electrostatic and flux bias [krantz2019quantum].
 For the CR gate, $\mathrm{Q_{d}}$ is excited by driving $\mathrm{Q_{b}}$ at the frequency of $\mathrm{Q_{d}}$, generating the transition $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$ (refs.
-[18, 19, 20]).
+[rigetti2010fully, magesan2020effective, chow2011simple]).
 For the bSWAP gate, $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ are driven monochromatically through a two-photon process near the frequency middle point between the two, generating the transition $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$ (refs.
-[17, 21]).
+[poletto2012entanglement, roth2017analysis]).
 Two-qubit spectroscopy
-We have achieved the strong coupling between the two eNe qubits.
-Figure Coherent manipulation of interacting electron qubits on solid neonA shows the two-tone qubit spectroscopy measurement.
-We drive the system with a continuous-wave (CW) tone at a variable frequency, and probe the system with another CW tone at the fixed resonator frequency, with a low power, approximately 130-130 dBm reaching the resonator’s input coupler.
-Two clear avoided crossings are observed while tuning the bias voltage on the trap guard electrode ($\Delta V_{\rm tg}$).
-The splitting magnitude of $J/2\pi=$ 3.35 MHz is larger than the linewidth of the hybrid states above and below the splitting, as shown in Fig.
+We have experimentally achieved strong charge-charge coupling between eNe qubits, assisted by the neon surface, as proposed.
+See Methods for the details of the superconducting microwave resonator, neon growth and electrons deposition.
+A two-qubit system was probed by the continuous wave (CW) two-tone qubit spectroscopy while tuning one electrode gate ($\Delta V_{\rm tg}$, see Methods).
+Two clear avoided crossings which indicate the qubit-qubit strong coupling appear in the spectrum in Fig.
+Coherent manipulation of interacting electron qubits on solid neona.
+One of the qubits is visible in the CW spectroscopy and referred to as the “bright” qubit $\mathrm{Q_{b}}$, while the other non-detectable one is denoted as the “dark” qubit $\mathrm{Q_{d}}$.
+The magnitude of the splitting represents an inter-qubit coupling strength of $J/2\pi$ = 3.35 MHz, exceeding the linewidth of the hybrid states $(\gamma_{\textrm{b}}+\gamma_{\textrm{d}})/4\pi$ = 0.56 MHz, as shown in Fig.
 Coherent manipulation of interacting electron qubits on solid neonb.
-This indicates the strong coupling between the “bright” qubit ($\mathrm{Q_{b}}$) and “dark” qubit ($\mathrm{Q_{d}}$).
-Individually, the bright $\mathrm{Q_{b}}$ has a hyperbolic frequency dependence on $\Delta V_{\rm tg}$, with a charge sweet spot (SS) located near 5.71 GHz, which is about 40 MHz above the resonator frequency at 5.668 GHz.
-The coupling between $\mathrm{Q_{b}}$ and the resonator dispersively pushes the resonator spectrum, especially when $\mathrm{Q_{b}}$ is tuned to its SS frequency (fig.
-S0.7).
-Based on this, we estimate the coupling strength between $\mathrm{Q_{b}}$ and the resonator to be $g_{\mathrm{b}}/2\pi=3.76$ MHz.
-In contrast, the interaction between the dark $\mathrm{Q_{d}}$ and the superconducting resonator is much weaker than that of $\mathrm{Q_{b}}$.
-Meanwhile, the secondary shift of the resonator caused by the dispersive push of $\mathrm{Q_{d}}$ on $\mathrm{Q_{b}}$ is also small.
-Together, these lead to a negligible $\mathrm{Q_{d}}$-state-dependent shift of the resonator frequency and the absence of visible $\mathrm{Q_{d}}$ spectroscopy in Fig.
-Coherent manipulation of interacting electron qubits on solid neonA.
-In addition, the spectroscopic symmetry of the two avoided crossing in against $\Delta V_{\text{tg}}$ indicates that the frequency of the dark $\mathrm{Q_{d}}$ is much less sensitive to $\Delta V_{\rm tg}$ compared to that for the bright $\mathrm{Q_{b}}$, with a nearly constant frequency of 5.726 GHz in the voltage scanning range of Fig.
-Coherent manipulation of interacting electron qubits on solid neonA.
-These observations all suggest the misalignment between $\mathrm{Q_{d}}$’s dipole and the microwave and direct current (DC) gating fields.
-We attribute the two-qubit coupling to charge-charge interactions at short distances.
-The spectroscopic observations reveal a coupled two-qubit system represented in Fig.
-Coherent manipulation of interacting electron qubits on solid neonB, with a finite $g_{\mathrm{b}}$ but negligible $g_{\mathrm{d}}$ in Eq.
-4.
-Due to the distinct coupling strengths of the two qubits individually with the resonator, we can rule out the case of virtual microwave photon exchange between the two qubits through the resonator, in which the inter-qubit exchange strength $J\,=\,g_{\mathrm{b}}g_{\mathrm{d}}(1/\Delta_{\mathrm{b}}+1/\Delta_{\mathrm{d}})/2$, where $\Delta_{\mathrm{b,d}}$ are the detune between the qubits and the resonator [22].
-The dark $\mathrm{Q_{d}}$ can be probed by swapping its population to the bright $\mathrm{Q_{b}}$.
-The swap operation can be realized using the ac-Stark shift of $\mathrm{Q_{b}}$ when the resonator is populated while probing the system with a readout pulse at the resonator frequency.
-Given the dispersive coupling strength of $\chi/2\pi\approx 0.33$ MHz, the ac-Stark effect will blue shift $\mathrm{Q_{b}}$’s frequency by $2\bar{n}\chi$ (refs.
-[6, 2]), crossing $\mathrm{Q_{d}}$’s frequency when the averaged intra-cavity photon number $\bar{n}$ is approximately 20.
-Due to the undetectable coupling between $\mathrm{Q_{d}}$ and the resonator, the probe signal will not affect $\mathrm{Q_{d}}$ directly.
-With the increase of photon number in the cavity during the readout with sufficient probe power, $\mathrm{Q_{b}}$’s frequency gradually crosses that of $\mathrm{Q_{d}}$, resulting in the states swap and eventually the $\mathrm{Q_{d}}$-state-dependent readout.
-This effect is turned off when the probe tone is weak, and the ac-Stark shift of $\mathrm{Q_{b}}$ is insufficient to cross the frequency of $\mathrm{Q_{d}}$, as in the case of Fig.
-Coherent manipulation of interacting electron qubits on solid neonA.
-We simulate the pulsed readout process in ref.
-[23], with varied probe power to illustrate the detailed swap process during a 700 ns readout pulse (fig.
-S0.7), as in the following experiments.
+Individually, $\mathrm{Q_{b}}$ has a hyperbolic frequency dependence on $\Delta V_{\rm tg}$, with a charge sweet spot (SS) near 5.71 GHz, which is about 40 MHz above the resonator frequency.
+We estimate the coupling strength between $\mathrm{Q_{b}}$ and resonator to be $g_{\mathrm{b}}/2\pi$ = 3.76 MHz, based on the dispersive pushing at the qubit’s SS (Extended Data Fig.
+1).
 Fig.
@@
 : Time-domain all-microwave two-qubit operations on solid neon.
-(A) and (C) Driving pulse length-frequency (A) and driving pulse amplitude-frequency (C) Rabi measurements, following the pulse sequences in the inset, showing the two operations corresponding to cross-resonance (CR, green arrows) and bSWAP (magenta arrows) two-qubit gates.
-(B) and (D) Simulated $\mathrm{Q_{b}}$ excited state population after the readout pulse in driving pulse length-frequency (B) and driving pulse amplitude-frequency (D) Rabi measurements.
-With this method, we perform pulsed qubit spectroscopy when biased at $\mathrm{Q_{b}}$’s SS.
-A long square pulse (10 µs) with a varying driving frequency $f_{\mathrm{dr}}$ and a driving power ($P_{\mathrm{dr}}$) is first applied to drive the system to a specific state, followed by a short square probe pulse (700 ns) at the resonator frequency $f_{\mathrm{r}}$ with approximately 120-120 dBm power reaching the resonator.
-The pulse sequence is shown in the inset of Fig.
+a and c, Drive pulse length-frequency (a) and drive pulse amplitude-frequency (c) Rabi measurements, following the pulse sequences depicted in the inset.
+In (a), the drive pulse power is fixed at 70-70 dBm, with varying length from 0 to 2.0 µs.
+In (c), the drive pulse length is fixed at 0.8 µs, with output amplitude varying from 0.0 V to 0.25 V, corresponding to zero power to 57-57 dBm reaching the resonator input coupler.
+The parameters of the readout pulses are the same as described in the Two-qubit Spectroscopy section.
+The system is biased with a single DC gate at $\Delta V_{\mathrm{tg}}=0$.
+Measurement results show the two oscillations corresponding to cross-resonance (CR, purple arrows) and bSWAP (magenta arrows) two-qubit gates.
+b and d, Numerically simulated excited state population of $\mathrm{Q_{b}}$ after the readout pulses, corresponding to the Rabi measurements in (b) and (d).
+The CR and bSWAP oscillation features are reproduced with experimentally measured parameters from the spectroscopy and decoherence characterization, see Methods for details.
+In contrast, the direct interaction between $\mathrm{Q_{d}}$ and the resonator is not detectable.
+Meanwhile, the secondary frequency shift of the resonator caused by the dispersive push of $\mathrm{Q_{d}}$ on $\mathrm{Q_{b}}$ is also small.
+Together, these lead to a negligible $\mathrm{Q_{d}}$-state-dependent resonator response and the absence of visible $\mathrm{Q_{d}}$ spectrum in Fig.
+Coherent manipulation of interacting electron qubits on solid neona.
+In addition, the spectroscopic symmetry of the two avoided crossings in against $\Delta V_{\text{tg}}$ indicates that the frequency of the dark qubit $\mathrm{Q_{d}}$ is much less sensitive to $\Delta V_{\rm tg}$ compared to that of $\mathrm{Q_{b}}$, remaining nearly constant at 5.726 GHz over the voltage range scanned in Fig.
+Coherent manipulation of interacting electron qubits on solid neona.
+These observations reveal that the $\mathrm{Q_{d}}$’s dipole is nearly orthogonal to both the MW and DC gate fields, mainly defined by the neon-induced structural confinement.
+We attribute the observed two-qubit coupling to charge-charge interactions at short distances.
+Due to the distinct electron-photon coupling strengths of the two qubits individually with the resonator, we can rule out the case of virtual microwave photon exchange between the two qubits through the resonator [majer2007coupling].
+Although we cannot readout the dark qubit through the resonator directly, bringing $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ on resonance allows their populations to swap, which reveals the state of $\mathrm{Q_{d}}$ prior to that.
+This resonance condition is achieved via the ac-Stark shift induced by the readout pulse.
+Once resonant, the states exchange, and a subsequent measurement of $\mathrm{Q_{b}}$ reveals the information of the original $\mathrm{Q_{d}}$ state.
+A weak pulse provides an insufficient shift, preventing the swap and leaving $\mathrm{Q_{d}}$ unmeasured, as in the CW two-qubit spectroscopy of Fig.
+Coherent manipulation of interacting electron qubits on solid neona.
+We simulated the pulsed readout process with varied probe power, as detailed in Methods, to illustrate the state swap dynamics during the 0.7 µs readout pulse used in our experiments (Extended Data Fig.
+4).
+With this method, we perform pulsed qubit spectroscopy when biased at $\mathrm{Q_{b}}$’s SS, as shown in Fig.
 Coherent manipulation of interacting electron qubits on solid neonc.
-At low driving power, we observe two transitions at 5.711 GHz and 5.726 GHz, corresponding to the frequencies of $\mathrm{Q_{b}}$ at its SS and $\mathrm{Q_{d}}$.
-The appearance of $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$ transition indicates that the long driving pulse pumps $\mathrm{Q_{d}}$ to a certain population, corresponding to the CR transition of a two-qubit coupled system [20, 18], as shown in Fig.
+At low drive power of 73-73 dBm reaching the resonator, we observe two transitions at 5.711 GHz and 5.726 GHz, corresponding to the frequencies of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$.
+The appearance of $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|0_{\mathrm{b}}1_{\mathrm{d}}\rangle$ transition indicates that the long drive pulse pumps $\mathrm{Q_{d}}$ to a certain population, corresponding to the CR transition of a two-qubit coupled system [chow2011simple, rigetti2010fully], as depicted in Fig.
 Coherent manipulation of interacting electron qubits on solid neonc.
-With an increasing driving power, we observe another transition peak ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$) near the middle of the $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ frequencies of 5.7185 GHz.
+With increasing drive power, an additional transition peak ($|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$) emerges near the midpoint between the $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ frequencies, located at 5.718 GHz.
 This corresponds to the bSWAP type of two-qubit operation in Fig.
-Coherent manipulation of interacting electron qubits on solid neonC, involving two-photon excitation [17, 21].
-The fact that $\omega_{|1_{\mathrm{b}}1_{\mathrm{d}}\rangle}-\omega_{|0_{\mathrm{b}}1_{\mathrm{d}}\rangle}-\omega_{|1_{\mathrm{b}}0_{\mathrm{d}}\rangle}\simeq 0$ suggests the effective longitudinal (ZZ) interaction in the system is negligible [11, 15].
+Coherent manipulation of interacting electron qubits on solid neonc, involving two-photon excitation [poletto2012entanglement, roth2017analysis].
+The fact that $2\omega_{|1_{\mathrm{b}}1_{\mathrm{d}}\rangle}\simeq\omega_{|0_{\mathrm{b}}1_{\mathrm{d}}\rangle}+\omega_{|1_{\mathrm{b}}0_{\mathrm{d}}\rangle}$ suggests the effective longitudinal (ZZ) interaction in the system is negligible [beysengulov2024coulomb, fors2024comprehensive].
 In summary, through the frequency-domain characterization, we have observed the coherent coupling between two qubits on solid neon, which can support both the CR and the bSWAP types of two-qubit operation.
 Two-qubit gate operations
-Following the frequency-domain characterization of the coupled system, we perform time-domain two-qubit operations corresponding to the CR and bSWAP gates.
-Figure Coherent manipulation of interacting electron qubits on solid neonA plots the readout signal following Gaussian-shaped driving pulses with varying lengths and approximately 70-70 dBm reaching the resonator input coupler.
-The driving pulses are truncated at $\pm 2.5\sigma$ on each side, where $\sigma$ is the standard deviation of the Gaussian envelope.
-The parameters of the readout pulses are the same as described in the last section.
-Two oscillation features corresponding to the CR operation near 5.726 GHz and the bSWAP operation near 5.718 GHz can be observed.
-Based on the system Hamiltonian, we numerically simulate the system evolution through the drive and readout process in the Rabi measurements [23, 24].
-Figure Coherent manipulation of interacting electron qubits on solid neonB plots the simulated population of $\mathrm{Q}_{\mathrm{b}}$ after readout.
-And the evolution of the two-qubit system through the driving process is shown in fig.
-S0.7.
-Thanks to the readout process, the excitation of $\mathrm{Q_{d}}$ via the two-qubit operation can be reflected in the $\mathrm{Q_{b}}$ population, as shown in Fig.
-Coherent manipulation of interacting electron qubits on solid neonB, and consequently, the recorded readout signal.
 Fig.
@@
 : Spectroscopic characterization of a three-qubit coupled system on solid neon.
-(A) Two-tone measured qubits spectroscopy showing the three-qubit coupled system.
+a, Two-tone qubit spectroscopy reveals a three-qubit coupled system, with the inter-qubit coupling strengths $J_{\mathrm{12}}$ and $J_{\mathrm{23}}$ shown in the inset.
+Via $\mathrm{Q}_{1}$, the system is probed at the resonator frequency while being driven with a second tone at varied frequencies.
 The dashed line marks the avoided crossing as in (b).
-(B) The line cut showing large coupling strength between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$ of $J_{\mathrm{12}}/2\pi=$ 62.5 MHz.
-(C) Calculated eigenstate energy diagram of the three-qubit coupled system, with bare qubit states shown as dashed lines.
+b, The line cut shows the large charge-charge interaction strength between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$ of $J_{\mathrm{12}}/2\pi=62.5$ MHz and hybrid linewidth $(\gamma_{1}+\gamma_{2})/4\pi=3.5$ MHz.
+c, Calculated eigenstate energy diagram of the three-qubit coupled system, with bare qubit states shown as dashed lines.
+See Methods for detailed qubit parameters used in the calculation.
+Following the frequency-domain characterization of the coupled system, we perform time-domain two-qubit operations corresponding to the CR and bSWAP gates.
+A Gaussian-shaped microwave pulse, with variable frequency and duration, excites the system and is followed by a readout pulse at the resonator frequency.
+This readout pulse also functions as the $\mathrm{Q_{b}}$ frequency tuning knob with ac-Stark effect and induces population exchange between $\mathrm{Q_{d}}$ and $\mathrm{Q_{b}}$.
+The CR and bSWAP gate operations are manifested as oscillations in the readout signal in Fig.
+Coherent manipulation of interacting electron qubits on solid neona, at 5.726 GHz and 5.718 GHz respectively.
+The simulation of this dynamic process with the system Hamiltonian gives post-readout population of $\mathrm{Q_{b}}$ with similar patterns, as shown in Fig.
+Coherent manipulation of interacting electron qubits on solid neonb.
+The detailed evolution of the two-qubit system through this drive-readout process is presented in Extended Data Fig.
+3.
 The simulated results match well with the experimental observations.
-For the CR operation (green arrows), $\mathrm{Q_{d}}$ was driven by pumping $\mathrm{Q_{b}}$ at the transition frequency of $\mathrm{Q_{d}}$, whose excited state population reached maximum when the pulse length was approximately 0.8 µs.
-Meanwhile, the CR operation barely excites $\mathrm{Q_{b}}$, as shown in fig.
-S0.7.
-For the bSWAP operation (magenta arrows), the two-photon process drives the two qubits simultaneously, while the oscillation frequency is much lower than the CR operation.
-Under a square driving pulse with amplitude $A$, and negligible direct driving on the $\mathrm{Q_{d}}$, the oscillation frequency induced by CR operation is $\Omega_{\text{CR}}=AJ/(\Delta_{\mathrm{bd}}+2J^{2}/\Delta_{\mathrm{bd}})$ (ref.
-[18]).
-For the bSWAP operation, the oscillation frequency is approximately $\Omega_{\text{bSWAP}}=2A^{2}J/(\Delta_{\mathrm{bd}}+2J^{2}/\Delta_{\mathrm{bd}})^{2}$ (ref.
-[17]), where $\Delta_{\mathrm{bd}}$ is the frequency separation between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$.
-When the bSWAP operation performs a $\pi$ rotation, which is about 2 µs in our case, it is locally equivalent to an iSWAP gate with appropriate single-qubit rotations [17].
-Furthermore, we performed driving pulse-amplitude-dependent Rabi oscillation measurement with a Gaussian-shaped driving pulse of 800 ns length and output pulse amplitude varying from 0.0 V to 0.25 V, corresponding to zero power to 57-57 dBm reaching the resonator input coupler.
-Again, we observe oscillating patterns near 5.726 GHz and 5.718 GHz, corresponding to the CR (green arrows) and bSWAP (magenta arrows) operations.
+For the CR operation (purple arrows), $\mathrm{Q_{d}}$ was driven by pumping $\mathrm{Q_{b}}$ at the transition frequency of $\mathrm{Q_{d}}$, whose excited state population reached maximum when the pulse length was approximately 0.8 µs.
+Meanwhile, the CR operation barely excites $\mathrm{Q_{b}}$, as shown in Extended Data Fig.
+3.
+For the bSWAP operation (magenta arrows), the two-photon process drives the two qubits simultaneously, while the oscillation frequency is lower than for the CR operation.
+When it performs a $\pi$ rotation, which is about 2 µs in our case, the bSWAP gate is equivalent to an iSWAP gate with appropriate single-qubit operations [poletto2012entanglement].
+Furthermore, we performed amplitude-dependent Rabi oscillation measurement with a Gaussian-shaped drive pulse of 0.8 µs duration.
+Again, we observe oscillating patterns near 5.726 GHz and 5.718 GHz, corresponding to the CR (purple arrows) and bSWAP (magenta arrows) operations.
 We attribute the “waterfall-like” patterns in the frequency range 5.713 through 5.718 GHz in Fig.
-Coherent manipulation of interacting electron qubits on solid neonC to the effects of the Gaussian pulse shape and the ac-Stark shift of $\mathrm{Q_{b}}$ under higher driving power [23], which are reproduced by simulations in Fig.
-Coherent manipulation of interacting electron qubits on solid neonD and fig.
-S0.7.
-These results demonstrate, for the first time, time-domain two-qubit operations in a eNe qubit platform, offering the potential for multi-qubit quantum processors based on charge qubits on solid neon.
-Further optimization of the two-qubit gate and readout is crucial to achieving high-fidelity operations.
+Coherent manipulation of interacting electron qubits on solid neonc to the effects of the Gaussian shape of the drive pulse and the ac-Stark shift of $\mathrm{Q_{b}}$ under higher drive power, which are reproduced by simulations in Fig.
+Coherent manipulation of interacting electron qubits on solid neond and Extended Data Fig.
+3.
+These results demonstrate, for the first time, time-domain two-qubit operations in the eNe qubit platform.
+The neon surface confinement assisted the global all-microwave gate with simplified local DC bias, offering the potential for compact multi-qubit integration on solid neon.
+Further optimization of the two-qubit gates and readout is crucial to achieving high-fidelity operations.
 Three-qubit strong coupling
-We further realized a three-qubit system on solid neon with strong inter-qubit coupling.
-Figure Coherent manipulation of interacting electron qubits on solid neonA plots the two-tone measured qubit spectrum while varying the gate voltage on the resonator electrode ($\Delta V_{\rm res}$), revealing the energy diagram of this three-qubit system.
-One of them ($\mathrm{Q}_{1}$) is directly coupled with a superconducting resonator, while the other two ($\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$) are serially coupled to $\mathrm{Q}_{1}$.
-Compared to $\mathrm{Q}_{1}$, the interactions strengths of $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ with the resonator are weak.
-We do not observe the vacuum Rabi splitting between them and the resonator.
-Besides, the dispersive phase shift of the resonator induced by $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ is also small when they are not hybridized with $\mathrm{Q}_{1}$.
-In Fig.
-Coherent manipulation of interacting electron qubits on solid neonB, the line cut around $\Delta V_{\rm res}=0.74$ mV shows a coupling strength of $J_{\textrm{12}}/2\pi=62.5$ MHz between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$.
-On the other hand, the coupling strength $J_{\textrm{23}}/2\pi$ between $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ is much smaller, at approximately 5 MHz level.
-The difference in inter-qubit coupling strengths can be caused by the varied distance and dipole orientation between qubits.
-Besides, since $\mathrm{Q}_{\textrm{2}}$ and $\mathrm{Q}_{\textrm{3}}$ both weakly coupled with the resonator, probably due to unfavorable alignment between their dipole and the microwave field, we attribute the inter-qubit coupling in the three-qubit system to charge-charge interactions instead of virtual photon exchange.
-Figure Coherent manipulation of interacting electron qubits on solid neonC shows the calculated eigenstates of the system, which match well with the two-tone spectroscopy.
-$\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$’s SS are both near 6.13 GHz, while $\mathrm{Q}_{3}$’s SS is more detuned to approximately 5.91 GHz.
-All the three qubits are responsive to the bias voltage on the resonator electrode with various sensitivities.
-The parameters applied in the calculation is listed in tab.
-S0.7.
+We further realized a three-qubit system on solid neon.
+The experiment was performed during a separate cooldown using a different resonator from that used in the previous two-qubit system.
+In this three-qubit system, $\mathrm{Q}_{1}$ is directly strong coupled with the resonator, manifested by vacuum Rabi splitting when they are on resonance, while there is no detectable coupling between the other two qubits with the resonator.
+The energy diagram of the coupled system was probed via a CW two-tone qubit spectroscopy, while varying the gate voltage on a nearby resonator electrode ($\Delta V_{\rm res}$, see Methods).
+The interactions between qubits are revealed by the avoided crossings in the system spectrum, as shown in Fig.
+Coherent manipulation of interacting electron qubits on solid neona.
+The line cut along the bias point where $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$ are on resonance, enlarged in Fig.
+Coherent manipulation of interacting electron qubits on solid neonb, shows a hybridize-induced avoided crossing with a splitting of $J_{\textrm{12}}/2\pi$ = 62.5 MHz.
+In the same spirit, the coupling strength between $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ is approximately $J_{\textrm{23}}/2\pi$ = 5 MHz.
+The difference in inter-qubit coupling strengths arises from variations in the distance and dipole orientation between qubits.
+Given that the couplings between $\mathrm{Q}_{2}$ and $\mathrm{Q}_{3}$ with the resonator are not detectable, likely because of an unfavorable alignment between their dipoles and the microwave field, we attribute the inter-qubit strong coupling to direct charge-charge interactions instead of virtual photon exchange via the resonator.
+Figure Coherent manipulation of interacting electron qubits on solid neonc shows the calculated eigenstates of the system in response to the bias voltage applied on the resonator electrode, aligning with the experimental two-tone spectroscopy.
+The parameters applied in the calculation are listed in Extended Data Tab.
+Coherent manipulation of interacting electron qubits on solid neon.
 Due to the large coupling strength between $\mathrm{Q}_{1}$ and $\mathrm{Q}_{2}$, the state mixing of the three qubits results in the energy diagram of the coupled system, which deviates from the uncoupled qubit spectrum.
@@
 Coherent manipulation of interacting electron qubits on solid neona.
-These observations indicate the possibilities of simultaneously entangling more than two qubits in the eNe system that may be applied for small- to medium-scale quantum simulations [14, 13].
+These observations indicate the possibilities of simultaneously entangling more than two qubits in the eNe system that may be applied for small- to medium-scale quantum simulations [wang2023atomic, hensgens2017quantum].
 Discussion and outlook
 In this work, we have realized coherent manipulation of multiple electron qubits on solid neon mediated by charge-charge interactions.
-In a two-qubit system, frequency-domain spectroscopy shows strong inter-qubit coupling.
-While time-domain measurements demonstrate both cross-resonance and bSWAP two-qubit operations, paving the way for the optimization of high-fidelity two-qubit gates on solid neon.
-Further, a three-qubit system with over 60 MHz inter-qubit coupling strength shows the potential of building more complex systems with qubits on solid neon.
-Charge-charge interactions between qubits at short distances may enable new possibilities for quantum information processing on solid neon.
+In the two-qubit system, frequency-domain spectroscopy shows strong inter-qubit coupling, while time-domain measurements demonstrate both cross-resonance and bSWAP two-qubit operations, paving the way for the optimizing high-fidelity two-qubit gates on solid neon.
+The bright qubit in the system act as the mediator between the dark qubit and the control and readout circuitry, showcasing the engineering potential of neon structures for functioning and scaling eNe qubits with compacted integration and suppressed wiring overheads.
+Further, the three-qubit system with over 60 MHz inter-qubit coupling strength shows the potential of building more complex systems with qubits on solid neon.
+Charge-charge interactions between qubits at short distances can enable new possibilities for quantum information processing on solid neon.
 For the two systems studied in this work, qubits can acquire distinct coupling strengths with superconducting resonators.
-Utilizing such possibilities, qubits with dipole moments oriented normal to the resonator field can serve as quantum memories for nearby coupled qubits [25, 26], or act as a quantum mediator [27].
-The variation in qubit-resonator coupling strengths further reflects the complexity of electron trapping on solid neon films and points to the important role of surface profile in determining qubit properties [9, 10, 28, 8].
-Establishing coherent charge interactions between adjacent qubits on solid neon is a crucial step in advancing this emerging solid-state qubit platform.
-Similarly, strong inter-qubit coupling has been demonstrated in semiconductor qubit dots through various interaction mechanisms at short distance [29, 30, 31, 32, 33].
-However, a multi-qubit system purely relying on local interactions will pose challenges to integrated control and readout circuits at large qubit volume.
-Entanglement between distant qubits through superconducting cavity buses could ease these limits posed by local interactions [22, 34, 35, 36].
-High-impedance resonators made of high-kinetic inductance thin films [37, 38] and Josephson junction arrays [39, 40] have been applied to achieve coherent interactions between distant semiconductor qubits.
-Further improvements in the control of individual qubit properties above solid neon with refined film growth, electron trapping and gating mechanisms, as well as the development of a compatible high-impedance resonator [41, 42] would facilitate remote entanglements between distant qubits and hybridize it with other quantum platforms [26, 43].
-Looking ahead, the integration of short- and long-distance interactions between qubits on solid neon could pave the way for an on-chip quantum network.
+Leveraging such possibilities, qubits isolated from the resonator can serve as quantum memories for nearby coupled qubits [wendin2017quantum, xie2024high], or act as a quantum mediator [baart2017coherent] in a qubit array.
+Enabling these applications requires further investigation into engineering the electron’s structural confinement with solid neon to control qubit properties and their interactions.
+[…]
+Methods
+Devices and experiments
+The data presented in Fig.
+Coherent manipulation of interacting electron qubits on solid neon and Fig.
+Coherent manipulation of interacting electron qubits on solid neon were collected on a high-impedance titanium nitride (TiN) splitting superconducting resonator, the same as the one in ref.
+[li2025noise].
+A schematic of the electron trap on the resonator is shown in Extended Data Fig.
+1a.
+The resonator supports a differential mode at 5.668 GHz with a total linewidth of $\kappa/2\pi$ = 0.38 MHz.
+Electrodes with on-chip low-pass filters are connected to the resonator pins and trap guards surrounding the electron-trapping region, enabling the application of bias voltages to tune the qubit transition frequency.
+The device was mounted on a home-made printed circuit board (PCB), sealed in a copper cell with neon filling lines, tungsten filaments, and electrical connectors on top of the cell lid.
+The cryogenic and room temperature electronics setups are the same as those in ref.
+[li2025noise].
+The data presented in Fig.
+Coherent manipulation of interacting electron qubits on solid neon were collected on a niobium splitting superconducting resonator [zhou2022single, zhou2024electron], which supports a differential mode at 6.426 GHz, with a total linewidth of $\kappa/2\pi$ = 0.46 MHz.
+The sample, cryogenic and room temperature electronic setups were the same as those in refs.
+[zhou2022single, zhou2024electron].
+The neon growth procedure is the same as that described in ref.
+[li2025noise].
+In summary, the dilution fridge is warmed up to create a temperature gradient from $\sim$27 K to $\sim$25 K between its 4K plate and mixing chamber (MXC) plate.
+At that moment, neon gas is supplied from a room-temperature gas handling system in the form of puffs and filled onto the device chip in the form of liquid.
+After a desired amount of neon filling, the fridge heater is turned off to let the fridge cool down to the base temperature.
+Electrons are deposited onto the chip at the base temperature by applying pulsed current to the tungsten filaments mounted above the device.
+During the electron firing process, temperature sensed on the MXC plate can increase up to 100 mK.
+The details of the procedure and discussion of the electron deposition process can be found in ref.
+[li2025noise].
+Modeling multi-qubit coupled systems
+The multi-qubit system Hamiltonian can be written as
+$\hat{H}_{\mathrm{sys}}=\hat{H}_{\mathrm{r}}+\hat{H}_{\mathrm{q}}+\hat{H}_{\mathrm{r\text{-}q}}+\hat{H}_{\mathrm{q\text{-}q}},$ (M1)
+where the resonator Hamiltonian $\hat{H}_{\mathrm{r}}$, the qubit Hamiltonian $\hat{H}_{\mathrm{q}}$, the resonator-qubit interaction Hamiltonian $\hat{H}_{\mathrm{r\text{-}q}}$, and the qubit-qubit direct interaction Hamiltonian $\hat{H}_{\mathrm{q\text{-}q}}$ are
+$\hat{H}_{\mathrm{r}}=\hbar\omega_{\mathrm{r}}\hat{a}^{\dagger}\hat{a}$, (M2)
+$\hat{H}_{\mathrm{q}}=\sum_{{i}}\frac{1}{2}\hbar\omega_{i}\hat{\sigma}_{{i}}^{\mathrm{z}},$ (M3)
+$\hat{H}_{\mathrm{r\text{-}q}}=\sum_{{i}}\hbar g_{i}(\hat{a}^{\dagger}\hat{\sigma}_{i}^{-}+\hat{a}\hat{\sigma}_{{i}}^{+}),$ (M4)
+and
+$\hat{H}_{\mathrm{q\text{-}q}}=\sum_{ij}\hbar J_{ij}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}+\hat{\sigma}_{i}^{-}\hat{\sigma}_{j}^{+})$ (M5)
+Here, $\omega_{\mathrm{r}}$ is the resonator frequency, $\omega_{i}$ is the transition frequency of $\mathrm{Q}_{i}$, $g_{i}$ is the coupling strength between the resonator and $\mathrm{Q}_{i}$, $\hat{a}$ and $\hat{a}^{\dagger}$ are the annihilation and creation operators of resonator photons, $\hat{\sigma}_{i}^{z}$, $\hat{\sigma}_{i}^{-}$, and $\hat{\sigma}_{i}^{+}$ are the Pauli-$z$, lowering, and raising operators acting on $\mathrm{Q}_{i}$.
+For the dark qubit, we have $g_{\mathrm{d}}\sim 0$.
+On the other hand, the bright qubit has a finite $g_{\mathrm{b}}$.
+$J_{ij}$ represents the transverse inter-qubit coupling strength between $\mathrm{Q}_{i}$ and $\mathrm{Q}_{j}$ (ref.
+[krantz2019quantum, poletto2012entanglement]).
+In the Two-qubit Spectroscopy section, we show that the experimentally observed longitudinal (ZZ) coupling strength is small in the system, so we exclude it from the system Hamiltonian model [fors2024comprehensive, beysengulov2024coulomb].
+The populations of qubits in coupled systems are modeled with their Hamiltonian and simulated using the QuTiP package [johansson2012qutip].
+For the numerical data presented in Fig.
+Coherent manipulation of interacting electron qubits on solid neon, we use Eq.
+M1 to Eq.
+M5 as the Hamiltonian, with parameters listed in the following sections and Extended Data Tab.
+Coherent manipulation of interacting electron qubits on solid neon.
+The Lindblad master equation is applied to model the effects of relaxation and dephasing:
+$\begin{split}\frac{d\hat{\rho}}{dt}\,=\,&-\frac{i}{\hbar}[\hat{H},\hat{\rho}]\\ &+\frac{\kappa}{2}(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\hat{a}^{\dagger}\hat{a}\hat{\rho}-\hat{\rho}\hat{a}^{\dagger}\hat{a})\\ &+\sum_{i}\frac{\Gamma_{i}}{2}(2\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})\\ &+\sum_{i}\frac{\Gamma_{i}^{\varphi}}{2}(2\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}-(\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})^{2}\hat{\rho}-\hat{\rho}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})^{2}),\end{split}$ (M6)
+where $\hat{H}=\hat{H}_{\mathrm{sys}}+\hat{H}_{\mathrm{dr}}$.
+$\hat{\rho}$ is the system density matrix.
+$1/\Gamma_{i}$ is the relaxation lifetime $T_{1}$ of $\mathrm{Q}_{i}$, and $2/\Gamma_{i}^{\varphi}$ is the pure dephasing time $T_{\varphi}$ of $\mathrm{Q}_{i}$, which are adopted from relaxation and Ramsey measurements in Extended Data Fig.
+2.
+To model the system under drive pulses starting at time $t=0$ with a time-dependent amplitude $A(t)$ and frequency $\omega_{\mathrm{dr}}$, the driving Hamiltonian can be written as
+$\hat{H}_{\mathrm{dr}}\,=\,A(t)\mathrm{cos}(t\omega_{\mathrm{dr}})\sum_{i}\hbar\eta_{i}(\hat{\sigma}_{i}^{+}+\hat{\sigma}_{i}^{-}),$ (M7)
+where $\eta_{i}\propto g_{i}$ is the ratio for direct drive on $\mathrm{Q}_{i}$ (ref.
+[poletto2012entanglement]).
+In the two-qubit system discussed in the main text $\eta_{\mathrm{b}}\sim 1$ and $\eta_{\mathrm{d}}\sim 0$.
+The numerical simulation is performed in the rotating frame at the drive frequency.
+To model the readout process, we introduce the Hamiltonian of the probing at the resonator ($\hat{H}_{\mathrm{p}}$) to the system Hamiltonian in the rotating frame as:
+[…]
+where $\epsilon$ is the probing amplitude.
+Details of the experimental characterizations of the system parameters, apart from the ones presented in the main text, are described in the following sections.
+Characterization of $\mathrm{Q_{b}}$-resonator coupling
+$\mathrm{Q_{b}}$ of the two-qubit coupled system presented in the main text is coupled to the resonator with its charge sweet-spot (SS) frequency above the resonator frequency.
+Due to the interaction, the resonator is dispersively shifted when $\mathrm{Q_{b}}$ is biased close to its SS, as shown in Extended Data Fig.
+1b.
+The line-cut of the resonator spectrum when ground-state $\mathrm{Q_{b}}$ is on its SS shows $\chi/2\pi=0.33$ MHz red shift compared to the bare resonator frequency, as shown in Extended Data Fig.
+1c.
+We estimated the coupling strength between $\mathrm{Q_{b}}$ and resonator to be $g_{\mathrm{b}}/2\pi\approx 3.76$ MHz based on this resonator dispersive shift and $\chi\approx g_{\mathrm{b}}^{2}/\Delta_{\mathrm{b}}$ (ref.
+[blais2021circuit]), where $\Delta_{\mathrm{b}}/2\pi=(\omega_{\mathrm{b}}-\omega_{\mathrm{r}})/2\pi\approx 43$ MHz.
+Characterization of qubits decoherence
+We evaluated $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$’s coherence via relaxation and Ramsey measurements when $\mathrm{Q_{b}}$ was biased at its SS.
+$\mathrm{Q_{b}}$ was directly driven by sending pumps through the resonator at its frequency.
+In contrast, $\mathrm{Q_{d}}$ was driven by pumping $\mathrm{Q_{b}}$ at $\mathrm{Q_{d}}$’s frequency, corresponding to the cross-resonance two-qubit operation.
+$\mathrm{Q_{d}}$ was probed with the pulsed readout method described in the main text and the previous sections.
+Extended Data Fig.
+2a and Extended Data Fig.
+2b show the relaxation of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ with $T_{1}$ of 1.8 µs and 30.3 µs, respectively, fitted from the measured in-phase signal.
+Extended Data Fig.
+2c shows the Ramsey fringes of $\mathrm{Q_{b}}$, with a $T_{2}^{*}$ of 2.56 µs fitted from the in-phase signal.
+The Ramsey measurement result of $\mathrm{Q_{d}}$ is more complicated, as we observed multiple frequency components of the measured in-phase signal, as shown in Extended Data Fig.
+2d and Extended Data Fig.
+2e.
+Numerical simulation of the Ramsey measurement process with a stable single frequency of $\mathrm{Q_{d}}$ reveals that the population swap between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ induced by the inter-qubit coupling will not affect the final probed state and measured signal.
+Therefore, we attribute the beating features in Extended Data Fig.
+2d to fluctuations of the $\mathrm{Q_{d}}$’s frequency during data acquisition, which has also occurred on other qubits measured on the same device, as presented in ref.
+[li2025noise].
+We fitted the data in Extended Data Fig.
+2d with a three-frequency model, which resulted in a $T_{2}^{*}$ of 32.99 µs.
+The measured relaxation and decoherence rates are then used in the numerical simulation of the system, as shown in Extended Data Tab.
+Coherent manipulation of interacting electron qubits on solid neon.
+Summary of simulation parameters
+Extended Data Tab.
+Coherent manipulation of interacting electron qubits on solid neon summarizes the simulation parameters to build the Hamiltonian of the two-qubit coupled system and to account for the decoherence.
+Since the longitudinal interaction is small, as discussed in the main text, ZZ coupling is excluded in the simulation.
+The resonator has a frequency of 5.668 GHz with linewidth $\kappa/2\pi=0.38$ MHz.
+The interaction strength between the bright qubit and the resonator is $g_{\mathrm{b}}/2\pi=3.76$ MHz.
+In addition, the drive amplitude $A$ in MHz used in the simulation is scaled with $A=735\times\mathcal{A}$, where $\mathcal{A}$ is the experimental drive pulse amplitude in $V$.
+The readout process is simulated with a 0.7 µs square pulse and probing amplitude $\epsilon$ at 18 MHz.
+Extended Data Fig.
+3 shows the simulated qubit excited state population before the readout process in the Rabi measurements, indicated as the arrow time stamp in the inset.
+The oscillation features at 5.726 GHz and 5.718 GHz, corresponding to the cross-resonance (CR) and bSWAP operation, match well with the experimental result in Fig.
+Coherent manipulation of interacting electron qubits on solid neon of the main text.
+It confirms that the observed oscillation features are the result of the two-qubit operations instead of the probing pulse following that.
+In Extended Data Fig.
+3a and Extended Data Fig.
+3c, when $\mathrm{Q_{d}}$ is excited by driving $\mathrm{Q_{b}}$ at 5.726 GHz, the $\mathrm{Q_{b}}$’s population barely changes.
+After the readout process, the $\mathrm{Q_{d}}$’s population is swapped onto $\mathrm{Q_{b}}$, as shown in Fig.
+Coherent manipulation of interacting electron qubits on solid neonb and Coherent manipulation of interacting electron qubits on solid neond of the main text.
+The CR and bSWAP oscillations have different frequencies.
+Assuming a square drive pulse with amplitude $A$, and negligible direct drive on the $\mathrm{Q_{d}}$, the oscillation frequency induced by CR operation is $\Omega_{\text{CR}}=AJ/(\Delta_{\mathrm{bd}}+2J^{2}/\Delta_{\mathrm{bd}})$ (ref.
+[rigetti2010fully]).
+For the bSWAP operation, the oscillation frequency is approximately $\Omega_{\text{bSWAP}}=2A^{2}J/(\Delta_{\mathrm{bd}}+2J^{2}/\Delta_{\mathrm{bd}})^{2}$ (ref.
+[poletto2012entanglement]), where $\Delta_{\mathrm{bd}}$ is the frequency separation between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$.
+To calculate the eigenstates of the three-qubit coupled system, as in the main text Fig.
+Coherent manipulation of interacting electron qubits on solid neon, we modeled qubits 1, 2, and 3 with hyperbolic frequency dependency on the gate voltage $\Delta V_{\rm res}$ as: $f_{i}=\sqrt{(\alpha_{i})^{2}+(\beta_{i}\times(\Delta V_{\rm res}-\delta_{i}))^{2}}$.
+In Extended Data Tab.
+Coherent manipulation of interacting electron qubits on solid neon, we list the parameters used to generate Fig.
+Coherent manipulation of interacting electron qubits on solid neonc in the main text.
+In addition, we used $J_{\textrm{12}}/2\pi\,=62.5$ MHz and $J_{\textrm{23}}/2\pi\,=5.0$ MHz.
+We also omit longitudinal interactions between qubits in this calculation.
+Pulsed readout of $\mathrm{Q_{d}}$
+We demonstrate the pulsed readout process of $\mathrm{Q_{d}}$ with the following simulation.
+Initially, the population of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ is set to be 0 and 1, respectively.
+Then the resonator is populated by sending in a square readout pulse (0.7 µs) at the resonator frequency with varying amplitude $\epsilon$.
+The evolution of the coupled system is calculated with the system Hamiltonian Eq.
+M1 and the probing Hamiltonian Eq.
+M8, considering relaxation and dephasing with Eq.
+M6.
+Extended Data Fig.
+4 shows the evolution of the intra-cavity photon number $\bar{n}$, $\mathrm{Q_{b}}$, and $\mathrm{Q_{d}}$ excited state population under varying drive amplitude.
+When the probe power is small (Extended Data Fig.
+4a and Extended Data Fig.
+4b), the ac-Stark shift of $\mathrm{Q_{b}}$ caused by the probe pulse is also small, $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ weakly exchange population at high frequency approximately $\sqrt{\Delta_{\mathrm{bd}}^{2}+4J^{2}}$, where $J$ is the inter-qubit coupling strength and $\Delta_{\mathrm{bd}}$ are the qubits detune.
+With increased photon number, the blue ac-Stark shift [blais2021circuit, zhou2024electron] $2\chi\bar{n}$ causes the $\mathrm{Q_{b}}$ across $\mathrm{Q_{d}}$ frequency.
+Given the dispersive coupling strength of $\chi/2\pi\approx 0.33$ MHz, the ac-Stark effect will result in the population swap when $\bar{n}$ approaches 20, as shown in Extended Data Fig.
+4c.
+Further increasing the probing amplitude could populate the resonator faster and make the population swap happen even sooner.
+These simulation results reveal the readout process we applied to probe $\mathrm{Q_{d}}$, whose direct interaction strength with the resonator is negligible.
+We chose the 0.7 µs readout length to balance signal strength with the fast decaying of $\mathrm{Q_{b}}$, as illustrated in the qubit decoherence section.
+Discussion of the “Waterfall” features in the pulse amplitude-dependent Rabi oscillation
+The “waterfall-like” pattern measured in pulse amplitude-dependent Rabi oscillation is also replicated in the numerical simulation, by realistically modeling the drive pulse with the same Gaussian envelope used in experiments.
+We attribute it to the combined effects of the pulse shape and ac-Stark shift of $\mathrm{Q_{b}}$ under drive at the frequency middle point between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ with high power.
+Extended Data Fig.
+5 plots the simulation when we replace the Gaussian-shaped drive pulse with square pulses.
+Under such conditions, the “waterfall-like” pattern vanishes while we could still observe the red shift of the $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$ transition at higher power, caused by the ac-Stark shift of $\mathrm{Q_{b}}$.
 Data availability
@@
 Schuster for helpful discussions.
-Supplementary Information for
-Coherent manipulation of interacting electron qubits on solid neon
-S0.1 Devices and experiments
-The data presented in Fig.
-Coherent manipulation of interacting electron qubits on solid neon and Fig.
-Coherent manipulation of interacting electron qubits on solid neon were collected on a high-impedance titanium nitride (TiN) splitting superconducting resonator, the same as the one in ref.
-[9].
-The resonator supports a differential mode at 5.668 GHz with a total linewidth of $\kappa/2\pi\,=\,$0.38 MHz.
-Electrodes with on-chip low-pass filters are connected to the resonator pins and trap guards surrounding the electron-trapping area to apply bias voltages for tuning qubits transition frequency.
-The device was mounted on a home-made printed circuit board (PCB), sealed in a copper cell with neon filling lines, tungsten filaments, and electrical connectors on top of the cell lid.
-The cryogenic and room temperature electronics setups are the same as those in ref.
-[9].
-The data presented in Fig.
-Coherent manipulation of interacting electron qubits on solid neon were collected on a niobium splitting superconducting resonator, the same as the one in refs.
-[1, 2].
-This resonator supports a differential mode at 6.426 GHz, with a total linewidth of $\kappa/2\pi=0.46$ MHz.
-The sample mounting, cryogenic and room temperature electronic setups were the same as those in refs.
-[1, 2].
-The neon growth procedure is the same as that described in ref.
-[9].
-In summary, the dilution fridge is warmed up to create a temperature gradient from $\sim$27 K to $\sim$25 K between its 4K plate and mixing chamber (MXC) plate.
-At that moment, neon gas is supplied from a room-temperature gas handling system in the form of puffs and filled onto the device chip in the form of liquid.
-After a desired amount of neon filling, the fridge heater is turned off to let the fridge cool down to base temperature.
-Electrons are deposited on to the chip by sending pulsed current on to the tungsten filaments mounted above the device chip.
-During the electron firing process, temperature sensed on the MXC plate can increase up to 100 mK.
-The details of the procedure and discussion of the electron deposition process can be found in ref.
-[9].
-S0.2 Simulation of multi-qubit coupled systems
-The populations of qubits in coupled systems are simulated based on their Hamiltonian, using the  QuTiP package [24].
-For the numerical data presented in Fig.
-Coherent manipulation of interacting electron qubits on solid neon, we use Eq.
-1 to Eq.
-5 as the Hamiltonian, with parameters listed in the following sections and Tab.
-S0.7.
-The Lindblad master equation is applied to model the effects of relaxation and dephasing:
-$\begin{split}\frac{d\hat{\rho}}{dt}\,=\,&-\frac{i}{\hbar}[\hat{H},\hat{\rho}]\\ &+\frac{\kappa}{2}(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\hat{a}^{\dagger}\hat{a}\hat{\rho}-\hat{\rho}\hat{a}^{\dagger}\hat{a})\\ &+\sum_{i}\frac{\Gamma_{i}}{2}(2\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})\\ &+\sum_{i}\frac{\Gamma_{i}^{\varphi}}{2}(2\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}\hat{\rho}\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}-(\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})^{2}\hat{\rho}-\hat{\rho}(\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-})^{2}),\end{split}$ (S1)
-where $\hat{H}=\hat{H}_{\mathrm{sys}}+\hat{H}_{\mathrm{dr}}$.
-$\hat{\rho}$ is the system density matrix.
-$1/\Gamma_{i}$ is the relaxation lifetime $T_{1}$ of $\mathrm{Q}_{i}$, and $2/\Gamma_{i}^{\varphi}$ is the pure dephasing lifetime $T_{\varphi}$ of $\mathrm{Q}_{i}$, which are adopted from relaxation and Ramsey measurements in Fig.
-S0.7.
-To simulate the system under driving pulses starting at time $t=0$ with a time-dependent amplitude $A(t)$ and frequency $\omega_{\mathrm{dr}}$, the driving Hamiltonian can be written as
-$\hat{H}_{\mathrm{dr}}\,=\,A(t)\mathrm{cos}(t\omega_{\mathrm{dr}})\sum_{i}\hbar\eta_{i}(\hat{\sigma}_{i}^{+}+\hat{\sigma}_{i}^{-}),$ (S2)
-where $\eta_{i}\propto g_{i}$ is the ratio for direct driving on $\mathrm{Q}_{i}$ (ref.
-[17]).
-In the two-qubit system discussed in the main text $\eta_{\mathrm{b}}\sim 1$ and $\eta_{\mathrm{d}}\sim 0$.
-The numerical simulation is performed in the rotating frame at the drive frequency.
-To simulate the readout process, we introduce the Hamiltonian of the probing at the resonator ($\hat{H}_{\mathrm{p}}$) to the system Hamiltonian in the rotating frame as:
-[…]
-where $\epsilon$ is the probing amplitude.
-S0.3 Resonator dispersive shift
-$\mathrm{Q_{b}}$ of the two-qubit coupled system presented in the main text is coupled to the resonator with its charge sweet-spot (SS) above the resonator frequency.
-Due to the interaction, the resonator is dispersively shifted when $\mathrm{Q_{b}}$ is biased close to its SS, as shown in Fig.
-S0.7A.
-The line-cut of the resonator spectrum when ground-state $\mathrm{Q_{b}}$ is on its SS shows $\chi/2\pi=0.33$ MHz red shift compared to the bare resonator frequency, as shown in Fig.
-S0.7B.
-We estimated the coupling strength between $\mathrm{Q_{b}}$ and resonator to be $g_{\mathrm{b}}/2\pi\approx 3.76$ MHz based on this resonator dispersive shift and $\chi\approx g_{\mathrm{b}}^{2}/\Delta_{\mathrm{b}}$ (ref.
-[6]), where $\Delta_{\mathrm{b}}/2\pi=(\omega_{\mathrm{b}}-\omega_{\mathrm{r}})/2\pi\approx 43$ MHz.
-S0.4 Pulsed readout of $\mathrm{Q_{d}}$
-We demonstrate the pulsed readout process of $\mathrm{Q_{d}}$ with the following simulation.
-Initially, the population of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ is set to be 0 and 1, respectively.
-Then the resonator is populated by sending in a square readout pulse (700 ns) at the resonator frequency with varying amplitude $\epsilon$.
-The evolution of the coupled system is calculated with the system Hamiltonian Eq.
-1 and the probing Hamiltonian Eq.
-S3, considering relaxation and dephasing with Eq.
-S1.
-Figure S0.7 shows the evolution of intra-cavity photon number $\bar{n}$, $\mathrm{Q_{b}}$, and $\mathrm{Q_{d}}$ excited state population under varying driving amplitude.
-When the probe power is small (Fig.
-S0.7A and Fig.S0.7B), the ac-Stark shift of $\mathrm{Q_{b}}$ caused by the probe pulse is also small, $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ weakly exchange population at high frequency approximately $\sqrt{\Delta_{\mathrm{bd}}^{2}+4J^{2}}$, where $J$ is the inter-qubit coupling strength and $\Delta_{\mathrm{bd}}$ are the qubits detune.
-With increased photon number, the blue ac-Stark shift $2\chi\bar{n}$ causes the $\mathrm{Q_{b}}$ across $\mathrm{Q_{d}}$ frequency, resulting in their population swap when $\bar{n}$ approaches 20, as shown in Fig.
-S0.7C.
-Further increasing the probing amplitude could populate the resonator faster and make the population swap happen even sooner.
-These simulation results reveal the readout process we applied to probe $\mathrm{Q_{d}}$, whose direct interaction strength with the resonator is negligible.
-We chose the 700 ns readout length to balance signal strength with the fast decaying of $\mathrm{Q_{b}}$, as illustrated in the qubit decoherence section.
-S0.5 Qubits decoherence
-We evaluated $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$’s coherence via relaxation and Ramsey measurements when $\mathrm{Q_{b}}$ was biased at its SS.
-$\mathrm{Q_{b}}$ was directly driven by sending pumps through the resonator at its frequency.
-In contrast, $\mathrm{Q_{d}}$ was driven by pumping $\mathrm{Q_{b}}$ at $\mathrm{Q_{d}}$’s frequency, corresponding to the cross-resonance two-qubit operation.
-$\mathrm{Q_{d}}$ was probed with the pulsed readout method described in the main text and the previous sections.
-Figure S0.7A and S0.7B show the relaxation of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ with $T_{1}$ of 1.8 µs and 30.3 µs, respectively, fitted from the measured in-phase signal.
-Figure S0.7C shows the Ramsey fringes of $\mathrm{Q_{b}}$, with a $T_{2}^{*}$ of 2.56 µs fitted from the in-phase signal.
-The Ramsey measurement result of $\mathrm{Q_{d}}$ is more complicated, as we observed multiple frequency components of the measured in-phase signal, as shown in Fig.
-S0.7D and  S0.7E.
-Numerical simulation of the Ramsey measurement process with a stable single frequency of $\mathrm{Q_{d}}$ reveals that the population swap between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ induced by the inter-qubit coupling will not affect the final probed state and measured signal.
-Therefore, we attribute the beating features in Fig.
-S0.7D to fluctuations of the $\mathrm{Q_{d}}$’s frequency during data acquisition, which has also occurred on other qubits measured on the same device, as presented in ref.
-[9].
-We fitted the data in Fig.
-S0.7D with a three-frequency model, which resulted in a $T_{2}^{*}$ of 32.99 µs.
-The measured relaxation and decoherence rates are then used in the numerical simulation of the system, as shown in Tab.
-S0.7.
-S0.6 Numerical simulation parameters
-Table S0.7 summarizes the simulation parameters to build the Hamiltonian of the two-qubit coupled system and to account for the decoherence.
-Since the longitudinal interaction is small, as discussed in the main text, ZZ coupling is excluded in the simulation.
-The resonator has a frequency of 5.668 GHz with linewidth $\kappa/2\pi=0.38$ MHz.
-The interaction strength between the bright qubit and the resonator is $g_{\mathrm{b}}/2\pi=3.76$ MHz.
-In addition, the driving amplitude $A$ in MHz used in the simulation is scaled with $A=735\times\mathcal{A}$, where $\mathcal{A}$ is the experimental driving pulse amplitude in $V$.
-The readout process is simulated with 700 ns square pulse with probing amplitude $\epsilon$ at 18 MHz.
-Figure S0.7 shows the simulated qubit excited state population before the readout process in the Rabi measurements, indicated as the arrow time stamp in the inset.
-The oscillation features at 5.726 GHz and 5.718 GHz, corresponding to the cross-resonance (CR) and bSWAP operation, match well with the experimental result in Fig.
-Coherent manipulation of interacting electron qubits on solid neon of the main text.
-It confirms that the observed oscillation features are the result of the two-qubit operations instead of the probing pulse following that.
-In Fig.
-S0.7A and S0.7C, when $\mathrm{Q_{d}}$ is excited by driving $\mathrm{Q_{b}}$ at 5.726 GHz, the $\mathrm{Q_{b}}$’s population barely changes.
-After the readout process, the $\mathrm{Q_{d}}$’s population is swapped onto $\mathrm{Q_{b}}$, as shown in Fig.
-Coherent manipulation of interacting electron qubits on solid neonB and Coherent manipulation of interacting electron qubits on solid neonD of the main text.
-To calculate the eigenstates of the three-qubit coupled system, as in the main text Fig.
-Coherent manipulation of interacting electron qubits on solid neon, we modeled qubits 1, 2, and 3 with hyperbolic frequency dependency on the gate voltage $\Delta V_{\rm res}$ as: $f_{i}=\sqrt{(\alpha_{i})^{2}+(\beta_{i}\times(\Delta V_{\rm res}-\delta_{i}))^{2}}$.
-In Tab.
-S0.7, we list the parameters used to generate Fig.
-Coherent manipulation of interacting electron qubits on solid neonC in the main text.
-In addition, we used $J_{\textrm{12}}/2\pi\,=62.5$ MHz and $J_{\textrm{23}}/2\pi\,=5.0$ MHz.
-We also omit longitudinal interactions between qubits in this calculation.
-S0.7 “Waterfall” features in pulse amplitude-dependent Rabi oscillation
-The “waterfall-like” pattern measured in pulse amplitude-dependent Rabi oscillation is also replicated in the numerical simulation, by realistically modeling the driving pulse with the same Gaussian envelope used in experiments.
-We attribute it to the combined effects of the pulse shape and ac-Stark shift of $\mathrm{Q_{b}}$ under driving at the frequency middle point between $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$ with high power.
-Figure S0.7 plots the simulation when we replace the Gaussian-shaped driving pulse with square pulses.
-Under such conditions, the “waterfall-like” pattern vanishes while we could still observe the red shift of the $|0_{\mathrm{b}}0_{\mathrm{d}}\rangle\rightarrow|1_{\mathrm{b}}1_{\mathrm{d}}\rangle$ transition at higher power, caused by the ac-Stark shift of $\mathrm{Q_{b}}$.
-Fig.
-S1.
-: Resonator dispersive shift caused by $\mathrm{Q_{b}}$.
-(A) Resonator spectrum probe with varying bias on trap guard, showing the dispersive shift when $\mathrm{Q_{b}}$ is close to resonator frequency.
-(B) Transmission line cuts at bare resonator (blue dots and curve) and when $\mathrm{Q_{b}}$ is biased at its sweet-spot (orange dots and curve).
-Fig.
-S2.
-: Simulation of intra-cavity photon number and qubit population evolution with varying probing amplitude.
-700 ns evolution simulated with probing amplitude $\epsilon$ of (A) 5.6 MHz (-130 dBm reaching the resonator), (B) 10 MHz (-125 dBm reaching the resonator) and (C) 18 MHz (-120 dBm reaching the resonator), respectively.
-Black curves show intra-cavity photon number $\bar{n}$.
-Blue curves show $\mathrm{Q_{b}}$ population and orange curves show $\mathrm{Q_{d}}$ population.
-Fig.
-S3.
-: Relaxation and decoherence of $\mathrm{Q_{b}}$ and $\mathrm{Q_{d}}$.
-(A) and (B) Relaxation measurements of $\mathrm{Q_{b}}$ (A) and $\mathrm{Q_{d}}$ (B) with fitted $T_{1}$ of 1.8 µs and 30.3 µs, respectively.
-(C) and (D) Ramsey measurements of $\mathrm{Q_{b}}$ (C) and $\mathrm{Q_{d}}$ (D) with fitted $T_{2}^{*}$ of 2.56 µs and 32.99 µs, respectively.
-(E) Frequency components of the Ramsey fringes in (D), revealing frequency fluctuations of $\mathrm{Q_{d}}$ during the measurement.
-(F) Numerical simulation of Ramsey measurement of $\mathrm{Q_{d}}$, indicating that the beating pattern in (C) is not intrinsically caused by the coupled system.
-Fig.
-S4.
-: Numerical simulation of qubit excited state population before the readout process in Rabi measurements.
-(A) and (B) Simulation results of $\mathrm{Q_{b}}$ (A) and $\mathrm{Q_{d}}$ (B) population after driving pulse in the pulse length-frequency Rabi measurements.
-(C) and (D) Simulation results of $\mathrm{Q_{b}}$ (C) and $\mathrm{Q_{d}}$ (D) population after driving pulse in the pulse amplitude-frequency Rabi measurements.
-Fig.
-S5.
-: Numerical simulation of pulse amplitude-dependent Rabi oscillation with step driving pulses.
-Simulation results of $\mathrm{Q_{b}}$ (A) and $\mathrm{Q_{d}}$ (B) population evolution, in which the “waterfall-like” pattern observed in experiments vanishes.
-Tab. S1. : Summary of the qubit parameters used for simulating the two-qubit coupled system as in Fig. Coherent manipulation of interacting electron qubits on solid neon and Fig. S0.7.
-$\omega_{i}/2\pi$ (GHz) $\Gamma_{i}/2\pi$ (MHz) $\Gamma_{i}^{\varphi}/2\pi$ (MHz) $g_{i}/2\pi$ (MHz) $J/2\pi$ (MHz)
-Bright $\mathrm{Q_{b}}$ 5.7112 0.088 0.036 3.76 3.35
-Dark $\mathrm{Q_{d}}$ 5.7255 0.0053 0.0044 0.0
-Tab. S2. : Summary of the qubit parameters used for modeling the three-qubit coupled system as in Fig. Coherent manipulation of interacting electron qubits on solid neonC of the main text.
-$\alpha_{i}$ (GHz) $\beta_{i}$ (GHz/mV) $\delta_{i}$ (mV)
-Qubit 1 6.130 1.325 0.0
-Qubit 2 6.135 0.218 -4.0
-Qubit 3 5.915 1.30 0.08